### v3

# Mining higher-order triadic interactions

Abstract
Complex systems often involve higher-order interactions which require us to go beyond their description in terms of pairwise networks. Triadic interactions are a fundamental type of higher-order interaction that occurs when one node regulates the interaction between two other nodes. Triadic interactions are found in a large variety of biological systems, from neuron-glia interactions to gene-regulation and ecosystems. However, triadic interactions have so far been mostly neglected. In this article, we propose a theoretical model that demonstrates that triadic interactions can modulate the mutual information between the dynamical state of two linked nodes. Leveraging this result, we propose the Triadic Interaction Mining (TRIM) algorithm to mine triadic interactions from node metadata, and we apply this framework to gene expression data, finding new candidates for triadic interactions relevant for Acute Myeloid Leukemia. Our work reveals important aspects of higher-order triadic interactions that are often ignored, yet can transform our understanding of complex systems and be applied to a large variety of systems ranging from biology to the climate.
**footnotetext: These authors contributed equally to this work
I Introduction
Higher-order networks [1, 2, 3, 4, 5] are key to capturing many-body interactions present in complex systems. Inferring higher-order interactions [6, 7, 8, 9, 10, 11] from real, pairwise network datasets is recognised as one of the central challenges in the study of higher-order networks [2, 12], with wide applicability across different scientific domains from biology and brain research [13, 14, 15] to finance [16, 17]. When from the knowledge of a pairwise network the problem is inferring the many-body interactions, mining higher-order interactions typically involves generative models and Bayesian approaches based on network structural properties [6, 18, 7, 9, 11]. Note, however that when the inference is performed on the basis on the knowledge of the nodes dynamical state [10, 8], inferring higher-order interactions also requires dynamical considerations.
Triadic interactions [19] are a fundamental type of signed higher-order interaction that are gaining increasing attention from the statistical mechanics community  [19, 20, 21, 22, 23, 24], since they are not reducible to hyperedges or simplices. A triadic interaction occurs when one or more nodes regulate the interaction between two other nodes. The regulator nodes may either enhance or inhibit the interaction between the other two nodes. Triadic interactions are known to be important in various systems, including: ecosystems [25, 26, 27] where one species can regulate the interaction between two other species; neuronal networks [28], where glial cells regulate synaptic transmission between neurons thereby regulating brain information processing; and gene regulation networks [29, 30], where a modulator can promote or inhibit the interaction between a transcription factor and a target gene. There is mounting evidence that triadic interactions can induce (so far largely unexplored) collective phenomena and dynamical states that can reveal important aspects of complex systems [19, 20, 21, 22, 23, 24, 31, 32]. An important advance in this line of research is triadic percolation [19, 20, 21], a theoretical framework that captures the non-trivial dynamics of the giant component. Moreover, recent results demonstrate that triadic interactions can have significant effects on stochastic dynamics [24] and learning  [22, 23]. However, despite the increasing attention that higher-order interactions are receiving, the detection of triadic interactions from network data and node timeseries, is an important scientific challenge that has not been thoroughly explored [33, 34, 29].
In this article, we formulate a dynamical model in which node variables are affected by triadic interactions, and we propose an information theoretic approach, leading to the Triadic Interaction Mining (TRIM) algorithm, for mining triadic interactions. The proposed dynamical model provides evidence of generative mechanisms by which a triadic interaction can induce a significant variability of the mutual information between two nodes at the end-points of an edge. The TRIM algorithm leverages this finding and mines triadic interactions using knowledge of the network structure and the dynamical variables associated to the nodes. The significance of each putative triadic interaction is then assessed using cross-validation with two distinct null models.
In this way, the TRIM algorithm can go beyond monotonicity assumptions regarding the functional form of the regulation of the two linked nodes by the third node (which is at the foundation of previously proposed methods [29]) allowing for broader applications. Significant node triples are also associated with an entropic score function $S$ that grows with the spread of the conditional joint distributions functions of the variables at the two end point of the regulated edge. We test the TRIM algorithm on a benchmark dynamical model, demonstrating its efficiency in detecting true triadic interactions. We also use the TRIM model to mine triadic interactions from gene-expression, to identify ‘trigenic’ processes [35]. We demonstrate that the TRIM algorithm is able to detect known interactions as well as propose a set of new candidate interactions that can then be validated experimentally.
Figure 1: (Panel a) A triadic interaction occurs when a node $Z$, called a regulator node, regulates (either positively or negatively) the interaction between two other nodes $X$ and $Y$. The regulated edge can be conceptualised as a factor node (shown here as a cyan diamond). (Panel b) A network with triadic interactions can be seen as a network of networks formed by a simple structural network and by a bipartite regulatory network between regulator nodes and regulated edges (factor nodes).
II Triadic interactions
A triadic interaction occurs when one or more nodes modulate (or regulate) the interaction between two other nodes, either positively or negatively. A triadic interaction network is a heterogeneous network composed of a structural network and a regulatory network encoding triadic interactions (Figure 1). The structural network $G_{S}=(V,E_{S})$ is formed by a set $V$ of $N$ nodes and a set $E_{S}$ of $L$ edges. The regulatory network $G_{R}=(V,E_{S},E_{R})$ is a signed bipartite network with one set of nodes given by $V$ (the nodes of the structural network), and another set of nodes given by $E_{S}$ (the edges of the structural network) connected by the regulatory interactions $E_{R}$ of cardinality $|E_{R}|=\hat{L}$. The signed regulatory network can be encoded as an ${L}\times N$ matrix $K$ where $K_{\ell i}=1$ if node $i$ activates the structural edge $\ell$, $K_{\ell i}=-1$ if node $i$ inhibits the structural edge $\ell$ and $K_{\ell i}=0$ otherwise. If $K_{\ell i}=1$, then the node $i$ is called a positive regulator of the edge $\ell$, and if $K_{\ell i}=-1$, then the node $i$ is called a negative regulator of the edge $\ell$. It is worth noting that node $i\in V$ cannot serve as both a positive and negative regulator for the same edge $\ell$ at the same time. However, node $i$ can act as a positive regulator for edge $\ell$ while simultaneously functioning as a negative regulator for a different edge $\ell^{\prime}\neq\ell$.
III The continuous triadic interaction model
[…]
$L_{ij}$ =$=\begin{cases}-a_{ij}J&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i\neq j,\\[2.5pt] \sum_{j}a_{ij}J&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i=j,\end{cases}$ (1)
where ${\bf A}=(a_{ij})$ is the adjacency matrix of the network, and $J>0$ is a coupling constant. In the absence of triadic interactions, we assume that the dynamics of the network is associated with a Gaussian process implemented as the Langevin equation
$d\mbox{\boldmath$X$}=-\frac{\delta\mathcal{H}}{\delta\mbox{\boldmath$X$}}+\Gamma\bm{\eta}(t),$ (2)
with the Hamiltonian
$\mathcal{H}=\frac{1}{2}{\mbox{\boldmath$X$}}^{\top}({\bf L}+\alpha{\bf I})\mbox{\boldmath$X$},$ (3)
where $\Gamma>0$, $\alpha>0$, and ${\bm{\eta}}(t)$ indicates uncorrelated Gaussian noise with
$\left\langle{\eta_{i}(t)\eta_{j}(t^{\prime})}\right\rangle=\delta_{ij}\delta(t-t^{\prime}),$ (4)
for all $t$ and $t^{\prime}$. The resulting Langevin dynamics are given by
$d\mbox{\boldmath$X$}=-(\mathbf{L}+\alpha\mathbf{I})\mbox{\boldmath$X$}dt+\Gamma d\bm{\eta}({t}).$ (5)
[…]
$C_{ij}=\frac{\Gamma^{2}}{2}[{\bf L}+\alpha{\bf I}]^{-1}_{ij}.$ (6)
In other words, from the correlation matrix it is possible to infer the Laplacian, and hence the connectivity of the network.
We now introduce triadic interactions in the model. As explained earlier, a triadic interaction occurs when one or more nodes modulate the interaction between another two nodes. To incorporate triadic interactions into the network dynamics, we modify the definition of the Laplacian operator present in the Langevin equation. Namely, we consider the Langevin dynamics
$d\mbox{\boldmath$X$}=-(\mathbf{L}^{(\text{T})}+\alpha\mathbf{I})\mbox{\boldmath$X$}dt+\Gamma d\bm{\eta}({t}),$ (7)
obtained from Eq.(2) by substituting the graph Laplacian ${\bf L}$ with the triadic Laplacian $\mathbf{L}^{(\text{T})}$ whose elements are given by
$L^{(\text{T})}_{ij}$ =$=\begin{cases}-a_{ij}J_{ij}(\mbox{\boldmath$X$})&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i\neq j,\\[2.5pt] \displaystyle\sum_{k=1}^{N}a_{ik}J_{ik}(\mbox{\boldmath$X$})&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i=j.\end{cases}$ (8)
Moreover, we assume that the coupling constants $J_{ij}(\mbox{\boldmath$X$})$ are determined by a perceptron-like model that considers all the regulatory nodes of the link $\ell=[i,j]$ and the sign of the regulatory interactions. Specifically, if $\sum_{k=1}^{N}K_{\ell k}X_{k}\geq\hat{T}$ then we set $J_{ij}=w^{+}$; if instead $\sum_{k=1}^{N}K_{\ell k}X_{k}<\hat{T}$ then we set $J_{ij}=w^{-}$, with $w_{+},w_{-}\in{}_{+}$ and $w_{+}>w_{-}$. Thus, we set
$J_{ij}(\mbox{\boldmath$X$})$ $=$ $w_{-}+(w_{+}-w_{-})\theta\left(\displaystyle\sum_{k=1}^{N}K_{\ell k}X_{k}-\hat{T}\right),$ (9)
[…]
Our model of triadic interactions, therefore, has the crucial property that triadic interactions are higher-order interactions that cannot be reduced to pairs of pairwise interactions. An important problem that then arises is whether such interactions can be mined from observational data. To address this issue, we will develop a new algorithm – that we call the TRIM algorithm – to identify triadic interactions from data, and we will test its performance on the data generated from the model described above.
Figure 2: The TRIM algorithm identifies triples of nodes $X$, $Y$, and $Z$ involved in a putative triadic interaction, starting from the knowledge of the structural network and the dynamical variables associated with its nodes. For each putative triple of nodes involved in a triadic interaction (panel (a)) which belong to a network whose structure and dynamics is known (panel (b)), we study the functional behavior of the conditional mutual information $MIz$ (panel (c)), and assess the significance of the observed modulations of $MIz$ with respect to a null model (panel (d)). Given a predetermined confidence level, we can use these statistics to identify significant triadic interactions (panel (e)). This procedure can be extended to different triples of the network, thereby identifying the triadic interactions present in it (panel (f)).
IV Mining triadic interactions
IV.1 The TRIM algorithm
We propose the TRIM algorithm (see Figure 2) to mine triadic interactions among triples of nodes. To simplify the notation we will use the letters $X,Y,Z$ to indicate both nodes as well as their corresponding dynamical variables.
Given a structural edge between nodes $X$ and $Y$, our goal is to determine a confidence level for the existence of a triadic interaction involving an edge between node $X$ and node $Y$ with respect to a potential regulator node $Z$. Specifically, we aim to determine whether the node $Z$ regulates the edge between node $X$ and node $Y$, given the dynamical variables $X$, $Y$ and $Z$ associated with these nodes. To do so, given a time series associated with node $Z$, we first sort the $Z$-values, and define $P$ bins in terms of the quantiles of $z$, chosen in such a way that each bin $m_{z}$ comprises the same number of data points (ranging in our analyses from $30$ to $100$). We indicate with $z_{m}$ the quantile of $Z$ corresponding to the percentile $m/P$. Therefore, each bin $m_{z}$ indicates data in which $Z$ ranges in the interval $[z_{m},z_{m+1})$. We indicate with $\mu(x|z_{m}),\mu(y|z_{m})$ and $\mu(x,y|z_{m})$ the probability density of the variables $X$, $Y$ and the joint probability density of the variables $X$ and $Y$ in each $m_{z}$ bin.
A triadic interaction is taken to occur when the node $Z$ affects the strength of the interaction between the other two nodes $X$ and $Y$. Consequently, our starting point is to consider the mutual information between the dynamical variables $X$ and $Y$ conditional to the specific value of the dynamical variable $Z$. We thus consider the quantity $MIz(m)=MI(X,Y|Z=z_{m})$ defined as
[…]
In order to estimate this quantity, we rely on non-parametric methods based on entropy estimation from $k$-nearest neighbors [41, 42, 43]. For each triple of nodes, we visualize the mutual information $MIz$ computed as a function of the $m/P$-th quantiles $z_{m}$ and fit this function with a decision tree comprising $r$ splits.
In the absence of triadic interactions, we expect $MIz$ to be approximately constant as a function of the $m/P$-th quantiles $z_{m}$, while, in the presence of triadic interactions, we expect this quantity to vary significantly as a function of $z_{m}$. The discretized conditional mutual information CMI between $X$ and $Y$ conditioned on $Z$ can be written as
$CMI_{X,Y;Z}=\sum_{m=0}^{P-1}p(z_{m})MIz(m)=\left\langle{MIz}\right\rangle,$ (10)
where $p(z_{m})=1/P$ indicates the probability that the $Z$ value falls in the $m_{z}$ bin. This quantity indicates important information about the interaction between the nodes $X$ and $Y$ when combined with the information coming from the mutual information MI given by
[…]
where $\mu(x),\mu(y)$ are the probability density functions for $X$ and $Y$ and $\mu(x,y)$ is the joint probability density functions of the variables $X$ and $Y$. In particular, true interactions between $X$ and $Y$ are typically associated with both high mutual information MI and high conditional mutual information CMI.
The conditional mutual information CMI, however, is not sensitive to variations in $MIz$ and does not therefore provide the information needed to detect triadic interactions. In order to overcome this limitation, we define the following three quantities that measure how much the mutual information between $X$ and $Y$ conditioned on $Z\in[z_{m},z_{m+1})$ changes as $z_{m}$ varies. Specifically we consider:
  • (1)
    the standard deviation $\Sigma$ of $MIz$, defined as
    $\Sigma=\sqrt{\sum_{m=0}^{P-1}p(z_{m}){\left[MIz(m)-\left\langle{MIz}\right\rangle\right]^{2}}};$ (11)
  • (2)
    the difference $T$ between the maximum and minimum value of $MIz$, given by
    $T=\max_{m=0,\ldots,P-1}\left|MIz(m)-\left\langle{MIz}\right\rangle\right|.$ (12)
The quantities $\Sigma$, $T$, collectively measure the strength of the triadic interaction under question and can thus be used to mine triadic interactions in synthetic as well as in real data. In order to assess the significance of the putative triadic interaction, we compare the observed values of these variables to the results obtained with given null models. In order to determine if the observed values are significant with respect to a given null model, we compute the scores $\Theta_{\Sigma}$, $\Theta_{T}$, given by
$\Theta_{\Sigma}$ $=$ $\frac{\Sigma-\mathbb{E}(\Sigma^{\textup{ran}})}{\sqrt{\mathbb{E}((\Sigma^{\textup{ran}})^{2})-(\mathbb{E}(\Sigma^{\textup{ran}}))^{2}}},$
$\Theta_{T}$ $=$ $\frac{T-\mathbb{E}(T^{\textup{ran}})}{\sqrt{\mathbb{E}((T^{\textup{ran}})^{2})-(\mathbb{E}(T^{\textup{ran}}))^{2}}},$
and the $p$-values
$p_{\Sigma}=\mathbb{P}(\Sigma^{\textup{ran}}>\Sigma),\quad p_{T}=\mathbb{P}(T^{\textup{ran}}>T),$ (14)
Note that if we consider $\mathcal{N}$ realizations of the null model, we cannot estimate probabilities smaller than $1/\mathcal{N}$. Therefore, if in our null model we observe no value of $\Sigma^{\textup{ran}}$ larger than the true data $\Sigma$, we set the conservative estimate $p_{\Sigma}=1/\mathcal{P}$. A similar procedure is applied also to $p_{T}$.
In this work we have considered two types of null models. The first is the randomization null model obtained by shuffling the $Z$ values, to give $\mathcal{N}$ randomization of the data, i.e. we use surrogate data for testing [44, 45]. The second is the maximum likelihood Gaussian null model between the three nodes involved in the triple $X$,$Y$,$Z$. We note that the adoption of these two null models allows us to also potentially capture non-monotonic relationships between $MIz$ and $z$, thereby going beyond underlying monotonic assumptions made in [46]. The first null model disrupts the temporal correlations between the timeseries of the node $Z$ and the timeseries the two nodes $X$ and $Y$ at the endpoints of the considered edges. Therefore this null model is robust with respect to the presence of possible outliers in the dataset. However this first null model may overlook confounding network effects that affect correlations between the dynamical variables. The second null model more efficiently captures correlations between the dynamical state of the three considered nodes due to network effects but is more sensitive to the presence of outliers in the data. To increase confidence, we therefore combine the insights coming from both null models (see SI for details).
For each triple, the function $MIz(m)$ is fitted with a decision tree with two splits. In this way, three different intervals of values of $Z$ are identified, each corresponding to a distinct functional behavior of the correlation functions between the variables $X$ and $Y$. While the method would allow for more than two splits of the decision three in general, here we have chosen three splits as this is the minimum number of splits that will allow us to capture non-trivial functional behaviour of $MIz$, like non-monotonicity. In practice this will be the best choice in presence of limited data and a small number of bins as in the Gene Expression data we will analyse in the following.
We further characterize significant triples by examining their entropic score function $S$, which is used to characterize their corresponding functional behaviour, as described in the SI.
As we will discuss below, the algorithm performs well on our paradigmatic model. Moreover, while the results obtained with the randomized surrogate data might neglect potentially important network effects, this shortcoming of this approach is mitigated by the cross-validation with the Gaussian null model.
Figure 3: Representative results for triples of nodes involved in triadic interactions in the continuous model with triadic interactions. The joint distributions of variables $X$ and $Y$ conditional on the values of $Z$ are shown in panel (a). Panel (b) shows the behavior of $MIz$ as a function of the values of $z_{m}$, which clearly departs from the constant behavior expected in absence of triadic interactions. Panel (c) presents the decision tree for fitting the $MIz$ functional behavior and determining the range of values of $Z$ for which the most significant differences among the joint distributions of the variables $X$ and $Y$ conditioned on $Z$ are observed. For more details about the simulation results, see the Supplementary Information (SI).
Figure 4: We consider a network with $N=10$, nodes $L=12$ edges, and $\hat{L}=5$ regulatory interactions (panel (a)). The time series obtained by integrating the stochastic dynamics of the proposed dynamical model for triadic interactions (Eq.(7)) are analyzed with the TRIM algorithm. Panel (b) displays the Receiver Operating Characteristic curve (ROC curve) obtained by running TRIM with $P=400$ bins and $\mathcal{N}=10^{3}$ realizations of the null model on these synthetic time series, using $\Theta_{\Sigma}$ to score for different parameters values indicated in the legend. The time series is simulated up to a maximum time $t_{max}=4000$ with a $dt=10^{-2}$. For the analysis, we consider $40000$ time steps (see the SI for details). The parameters of the model are: $\hat{T}=10^{-3}$, $w^{+}=8$, $w^{-}=0.5$, and $\alpha$ and $\Gamma$ as indicated in the figure legend.
IV.2 Validation of the TRIM algorithm on the triadic interaction model
To demonstrate the efficacy of the TRIM algorithm we first considered a single network (see Figure 4$(a)$) of $N=10$ nodes, $L=12$ edges and $\hat{L}=5$ triadic interactions (each formed by a single node regulating a single edge) on top of which we consider the triadic dynamical model proposed in Sec. III.
Simulations of Eq. (7) on this network show a strong dependence of $MIz(m)$ on $m$ for the triples of nodes involved in triadic interactions, with greater significance for smaller values of $\alpha$. Figure 3 shows the joint distributions $P_{\alpha}(X,Y)$ of $X$ and $Y$ for each interval $\alpha$ of values $Z$ determined by the decision tree. The results provide evidence of this interesting dynamical behavior of the triadic model in the case of a positive regulatory interaction. Note that our analysis of the function $MIz(m)$ also allows us to distinguish between positive and negative regulatory interactions. These are associated respectively by an increase in $MIz$ for larger values of $z_{m}$ or decrease of $MIz$ In the Supplementary Figures S1-S2, we display further examples of the function $MIz$ for triadic triples. We observe that as the parameter $\Gamma$ is raised, i.e. the noise increases, which is reflected in the increased variability of the $MIz$ functional behavior.
In order to provide an overall assessment of the performance of TRIM in mining triadic interactions on this network, in Figure 4$(b)$ we display the Receiver Operating Characteristic curve (ROC curve) based on the score $\Theta_{\Sigma}$, measured with respect to the first null model which we found to perform well for a wide range of parameters. As it is apparent from the ROC curve, the TRIM algorithm performs better for higher values of $\alpha$ and smaller values of $\Gamma$ (see Figure 4$(b)$). For any values of the parameters, we notice that the false positives are more likely to involve short-range triples, i.e. triples in which the regulator node $Z$ has short (small) structural network distance with the two nodes $X$ and $Y$.
Figure 5: Performance of the TRIM algorithm on a benchmark network with triadic interactions. Each data point shows measures for a triple of nodes $X$, $Y$ and $Z$, namely $\Theta_{\Sigma}$ and CMI, the proposed $\Sigma$-score and conditional mutual information between $X$ and $Y$, respectively. The colour of each point corresponds to the value of $S$, which characterises the entropy of the triple. The synthetic data comes from a network with $100$ nodes, $110$ edges, and $25$ triadic interactions. All possible triples are represented in the scatter plot. Stars are the true triadic triples.
We conclude our discussion by testing the TRIM algorithm on a much larger model network, to demonstrate its scalability. To this end, we consider a random network of $100$ nodes, $110$ edges, and $25$ triadic interactions (see SI for details). In Figure 5 we report the results of the TRIM algorithm for all the triples involving all edges, $14$ of which come from the pool of edges with triadic interactions. For each edge, all possible triples are analysed while for each edges we retain only the $5$ most significant triples according to $\Theta_{\Sigma}$. By conditioning on the value of third node, for each of these connected nodes we also record the conditional mutual information CMI. In figure 5, each considered triple corresponds to a point, colour coded according to the value of $S$. Stars indicate triples that are involved in a triadic interactions (see SI for details). The true triadic interactions are indeed found for triples with high $\Theta_{\Sigma}$ score, while CMI and $S$ span between high and intermediate values.
Figure 6: Maximal Spanning Tree of the relevant genes from the gene expression data. Edges and their edge weights are obtained from the Protein-Protein interaction network. Red-coloured nodes are nodes with biological significance, that is that play a critical role in AML[47][48][49]. Node size is proportional to the node degree.
Figure 7: (A) Scatter plot which displays $\Theta_{\Sigma}$ (Y-axis) and the CMI (X-axis) value between two genes, denoted $X$ and $Y$ , conditioned by a third gene, denoted $Z$. The colour corresponds to the value of our entropy measure $S$. Here only triples are depicted that rank among the 300 highest for CMI and MI and have $p$-value 0.001 or less in the randomization null model. Circles are triples resulting from the MST, whereas squares are triples involving genes with biological relevance. Crosses are triples that have $p$-value biggger than $0.0005$ with the Gaussian version null model. (B) The results of the decision tree allow us to distinguish three intervals in the value of Z in which the conditional distribution of the gene expression of $X$ and $Y$ differs the most. Plots of these conditional distributions are provided for two exemplary triples: both display a high significance score for the triadic interaction, indicating a meaningful biological association. The top row shows three biologically relevant genes, (X = HOXB3, Y = MEIS1, Z = GLIS3, $\Theta_{\Sigma}$ = 5.2, $p_{\Sigma}$ = 0.0001, $\Sigma$ = 0.38); the bottom row shows a triples identified by the MST. (C) has $p_{\Sigma}$-value 0.0001 in the random version. (X = GATA1, Y = KLF1, Z = ETV1, $\Theta_{\Sigma}$ = 6.1, $\Sigma$ = 0.44)
V Detecting triadic interactions in gene-expression data
Searching for trigenic interactions in gene-expression is a problem of major interest in biology. For instance, understanding the extent to which a modulator promotes or inhibits the interplay between a transcription factor and its target gene is crucial for deciphering gene regulation mechanisms. In order to address this question with our method, we considered gene-expression data from Acute Myeloid Leukemia (AML), extracted from the Grand Gene Regulatory Network Database [50, 51].
Exhaustive mining of all potential triadic interactions among each triple of nodes in the AML dataset is computationally very demanding (it would require testing of $>260M$ triplets). Instead, we here focus on edges associated with known biophysical interactions, as identified in the human Protein-Protein Interaction network (PPI)  [50]. To do so, we considered the subgraph of the human PPI network that contains all the genes/proteins included in the AML gene expression data. The number of edges in this subgraph is $M=42,511$. We first selected a set of triples involving genes known to be associated to AML with gene X and Y connected by an edge in the PPI (see SI for details). Moreover we selected additional triples according to a criteria based on the positions of the three genes in the Maximum Spanning Tree (MST) which only includes 621 edges (see Figure 6). For each edge in the MST, connecting gene X with gene Y we consider all genes Z within a distance of 4 from both the X and Y as candidate regulatory nodes, i.e. the third node in the triple (see SI for details). For each considered triple of genes we use $\Theta_{\Sigma}$ as the significance score, with $P=5$ bins, using $\mathcal{N}=5\times 10^{3}$ realizations of the randomization null model. These results have strong correlations with the results obtained considering $\Theta_{T}$ as the significance score (see Supplementary Figures S3-S4 for comparison).
Figure 7 shows the result of the TRIM algorithm for a selected set of the analysed triples with $p_{\Sigma}<0.001$. Specifically, the figure displays each triple in the plane $\Theta_{\Sigma}$ and CMI with a colour code encoding the value $S$. Note that for each selected edge only the top $5$ triples ranked according to the $\Theta_{\Sigma}$ score are depicted. Squares indicate triples chosen from biologically relevant genes for AML. Crosses indicate triples deemed insignificant in comparison with the Gaussian null model, Figures 7B and 7C provide examples of conditional distributions of two exemplary triples with high entropic score $S$. Both triples show evidence of triadic interactions: the triple in panel B is a member of the MST, while C is an example of a triple chosen from known biologically relevant genes for AML. For further examples of results obtained on more gene-expression triples, see the SI for details. Interestingly, among the significant triples, we detected also triples in which the modulation of the mutual information is non-monotonic (see SI for details).
Many of the genes involved in the $50$ highest ranking triples have already been linked to AML in the literature (see SI for Table 2 for a list of highly significant triples and Figure S6 with links to the literature associating the involved genes with AML). Interestingly, genes such as PAX4, SOX14 and VAX2 which appear in multiple triples have been found to be over- or under-expressed in patients with AML. Remarkably, 2 such genes, ASCL1 and NEUROG2, are involved in $5$ triples, suggesting that they may play a particularly important part in disease progression [52]. In total, $61\%$ of genes in the triples have a known association with AML.
VI Conclusions
This work provides a comprehensive information theory-based framework to model and mine triadic interactions directly from dynamic observations and for previous knowledge of the underlying pairwise network. The generative model we propose demonstrates that the presence of a triadic interaction leads to the modulation of the mutual information between the two end nodes of an edge (X and Y) by a third regulatory node Z. Via this model we show that to detect triadic interactions it is necessary to go beyond standard pairwise measures, such as the mutual information. Moreover, also standard higher-order statistical measures such as the conditional mutual information that only account for the average effect of the third regulatory node Z on the mutual information between X and Y falls short in characterizing triadic interactions. Our proposed approach, implemented in the TRIM algorithm, mines triadic interactions based on the quantification of the statistically significant modulation of the mutual information between the two linked nodes conditioned on the third regulator node. This algorithm is tested and validated on our generative model using paradigmatic random networks with triadic interactions and used to mine putative triadic interactions from gene expression data.
From the network theory point of view, this work opens new perspectives in the active field of modelling and inference of higher-order interactions and can be extended in many different directions, for instance by exploring the effect of triadic interactions on the dynamical state of nodes associated with discrete variables. From the biological point of view, our results may inspire further information-theoretic approaches to genetic regulatory network inference. Investigating the extent to which triadic interactions are tissue-specific, and if certain regulatory patterns are conserved across different tissues, could yield valuable insights. Our proposed approach could also be used to mine triadic interactions in other domains, such as finance or climate, where triadic interactions also have a significant role.
Code availability
The Python package triaction is available on GitHub at the following link: https://github.com/anthbapt/triaction
Acknowledgments
This research utilized Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT, https://doi.org/10.5281/zenodo.438045. M.N. A.B., R. S.-G. and G. B. acknowledge support from the Turing-Roche North Star partnership.

### v4

@@
 However, triadic interactions have so far been mostly neglected.
-In this article, we propose a theoretical model that demonstrates that triadic interactions can modulate the mutual information between the dynamical state of two linked nodes.
-Leveraging this result, we propose the Triadic Interaction Mining (TRIM) algorithm to mine triadic interactions from node metadata, and we apply this framework to gene expression data, finding new candidates for triadic interactions relevant for Acute Myeloid Leukemia.
-Our work reveals important aspects of higher-order triadic interactions that are often ignored, yet can transform our understanding of complex systems and be applied to a large variety of systems ranging from biology to the climate.
+In this article, we propose the Triadic Perceptron Model (TPM) that demonstrates that triadic interactions can modulate the mutual information between the dynamical state of two linked nodes.
+Leveraging this result, we formulate the Triadic Interaction Mining (TRIM) algorithm to extract triadic interactions from node metadata, and we apply this framework to gene expression data, finding new candidates for triadic interactions relevant for Acute Myeloid Leukemia.
+Our work reveals important aspects of higher-order triadic interactions that are often ignored, yet can transform our understanding of complex systems and be applied to a large variety of systems ranging from biology to climate.
 **footnotetext: These authors contributed equally to this work
@@
 Higher-order networks [1, 2, 3, 4, 5] are key to capturing many-body interactions present in complex systems.
-Inferring higher-order interactions [6, 7, 8, 9, 10, 11] from real, pairwise network datasets is recognised as one of the central challenges in the study of higher-order networks [2, 12], with wide applicability across different scientific domains from biology and brain research [13, 14, 15] to finance [16, 17].
-When from the knowledge of a pairwise network the problem is inferring the many-body interactions, mining higher-order interactions typically involves generative models and Bayesian approaches based on network structural properties [6, 18, 7, 9, 11].
-Note, however that when the inference is performed on the basis on the knowledge of the nodes dynamical state [10, 8], inferring higher-order interactions also requires dynamical considerations.
+Inferring higher-order interactions [6, 7, 8, 9, 10, 11] from real, pairwise network datasets is recognised as one of the central challenges in the study of higher-order networks [2, 12], with wide applicability across different scientific domains, from biology and brain research [13, 14, 15] to finance [16, 17].
+Mining higher-order interactions from the exclusive knowledge of the pairwise networks typically involves generative models and Bayesian approaches based on network structural properties [6, 18, 7, 9, 11].
+Note, however, that when the inference is performed on the basis of the knowledge of the nodes’ dynamical states [10, 8], inferring higher-order interactions also requires dynamical considerations.
 Triadic interactions [19] are a fundamental type of signed higher-order interaction that are gaining increasing attention from the statistical mechanics community [19, 20, 21, 22, 23, 24], since they are not reducible to hyperedges or simplices.
@@
 The regulator nodes may either enhance or inhibit the interaction between the other two nodes.
-Triadic interactions are known to be important in various systems, including: ecosystems [25, 26, 27] where one species can regulate the interaction between two other species; neuronal networks [28], where glial cells regulate synaptic transmission between neurons thereby regulating brain information processing; and gene regulation networks [29, 30], where a modulator can promote or inhibit the interaction between a transcription factor and a target gene.
-There is mounting evidence that triadic interactions can induce (so far largely unexplored) collective phenomena and dynamical states that can reveal important aspects of complex systems [19, 20, 21, 22, 23, 24, 31, 32].
+Triadic interactions are known to be important in various systems, including: ecosystems [25, 26, 27] where one species can regulate the interaction between two other species; neuronal networks [28], where glial cells regulate synaptic transmission between neurons thereby controlling brain information processing; and gene regulatory networks [29, 30], where a modulator can promote or inhibit the interaction between a transcription factor and its target gene.
+There is mounting evidence that triadic interactions can induce collective phenomena and/or modulate dynamical states that reveal important aspects of complex system behavior [19, 20, 21, 22, 23, 24, 31, 32].
 An important advance in this line of research is triadic percolation [19, 20, 21], a theoretical framework that captures the non-trivial dynamics of the giant component.
@@
 However, despite the increasing attention that higher-order interactions are receiving, the detection of triadic interactions from network data and node time series, is an important scientific challenge that has not been thoroughly explored [33, 34, 29].
-In this article, we formulate a dynamical model in which node variables are affected by triadic interactions, and we propose an information theoretic approach, leading to the Triadic Interaction Mining (TRIM) algorithm, for mining triadic interactions.
-The proposed dynamical model provides evidence of generative mechanisms by which a triadic interaction can induce a significant variability of the mutual information between two nodes at the end-points of an edge.
-The TRIM algorithm leverages this finding and mines triadic interactions using knowledge of the network structure and the dynamical variables associated to the nodes.
-The significance of each putative triadic interaction is then assessed using cross-validation with two distinct null models.
+In this article, we formulate the Triadic Perceptron Model (TPM) in which continuous node variables are affected by triadic interactions.
+Based on the insights gained by investigating this model, we propose an information theoretic approach, leading to the Triadic Interaction Mining (TRIM) algorithm, for mining triadic interactions.
+The TPM provides evidence of the mechanisms by which a triadic interaction can induce a significant variability of the mutual information between two nodes at the end-points of an edge.
+The TRIM algorithm leverages this finding and mines triadic interactions using knowledge of the network structure and the dynamical variables associated with the nodes.
+The significance of each putative triadic interaction is then validated by comparison with two distinct null models.
 In this way, the TRIM algorithm can go beyond monotonicity assumptions regarding the functional form of the regulation of the two linked nodes by the third node (which is at the foundation of previously proposed methods [29]) allowing for broader applications.
-Significant node triples are also associated with an entropic score function $S$ that grows with the spread of the conditional joint distributions functions of the variables at the two end point of the regulated edge.
-We test the TRIM algorithm on a benchmark dynamical model, demonstrating its efficiency in detecting true triadic interactions.
+Significant node triples are also associated with an normalized entropic score function $S\in[0,1]$ that quantifies the spread of the conditional joint distribution functions of the variables at the ends of the regulated edge.
+We test the TRIM algorithm on the benchmark TPM, demonstrating its efficiency in detecting true triadic interactions.
 We also use the TRIM model to mine triadic interactions from gene-expression, to identify ‘trigenic’ processes [35].
@@
 Figure 1: (Panel a) A triadic interaction occurs when a node $Z$, called a regulator node, regulates (either positively or negatively) the interaction between two other nodes $X$ and $Y$.
-The regulated edge can be conceptualised as a factor node (shown here as a cyan diamond).
+The regulated edge can be conceptualized as a factor node (shown here as a cyan diamond).
 (Panel b) A network with triadic interactions can be seen as a network of networks formed by a simple structural network and by a bipartite regulatory network between regulator nodes and regulated edges (factor nodes).
@@
 However, node $i$ can act as a positive regulator for edge $\ell$ while simultaneously functioning as a negative regulator for a different edge $\ell^{\prime}\neq\ell$.
-III The continuous triadic interaction model
-[…]
-$L_{ij}$ =$=\begin{cases}-a_{ij}J&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i\neq j,\\[2.5pt] \sum_{j}a_{ij}J&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i=j,\end{cases}$ (1)
-where ${\bf A}=(a_{ij})$ is the adjacency matrix of the network, and $J>0$ is a coupling constant.
+III The Triadic Perceptron Model (TPM)
+Here, we formulate a model for node dynamics in a network with triadic interactions that we call Triadic Perceptron Model (TPM).
+The TPM acts as a benchmark to validate the TRIM algorithm proposed here.
+We assume that each node $i$ of the network is associated with a dynamical variable $X_{i}\in\mathbb{R}$, and that the dynamical state of the entire network is encoded in the state vector $\mbox{$X$}=(X_{1},X_{2}\ldots,X_{N})^{T}$.
+The topology of the structural network is encoded in the graph Laplacian matrix $\bf L$} with elements
+$L_{ij}$ $=\begin{cases}-a_{ij}J&\text{if}\penalty 10000\ \penalty 10000\ i\neq j,\\[2.5pt] \sum_{k}a_{ik}J&\text{if}\penalty 10000\ \penalty 10000\ i=j,\end{cases}$ (1)
+where $\bf a$} is the adjacency matrix of the network of elements $a_{ij}$, and $J>0$ is a coupling constant.
 In the absence of triadic interactions, we assume that the dynamics of the network is associated with a Gaussian process implemented as the Langevin equation
-$d\mbox{\boldmath$X$}=-\frac{\delta\mathcal{H}}{\delta\mbox{\boldmath$X$}}+\Gamma\bm{\eta}(t),$ (2)
+$\frac{d\mbox{$X$}}{dt}=-\frac{\delta\mathcal{H}}{\delta\mbox{$X$}}+\Gamma\bm{\eta}(t),$ (2)
 with the Hamiltonian
-$\mathcal{H}=\frac{1}{2}{\mbox{\boldmath$X$}}^{\top}({\bf L}+\alpha{\bf I})\mbox{\boldmath$X$},$ (3)
-where $\Gamma>0$, $\alpha>0$, and ${\bm{\eta}}(t)$ indicates uncorrelated Gaussian noise with
-$\left\langle{\eta_{i}(t)\eta_{j}(t^{\prime})}\right\rangle=\delta_{ij}\delta(t-t^{\prime}),$ (4)
+$\mathcal{H}=\frac{1}{2}{\mbox{$X$}}^{\top}({\bf L}+\alpha{\bf I})\mbox{$X$},$ (3)
+where $\Gamma>0$, $\alpha>0$, and where ${\bm{\eta}}(t)$ indicates uncorrelated Gaussian noise with
+$\left\langle{\eta_{i}(t)}\right\rangle=0,\quad\left\langle{\eta_{i}(t)\eta_{j}(t^{\prime})}\right\rangle=\delta_{ij}\delta(t-t^{\prime}),$ (4)
 for all $t$ and $t^{\prime}$. The resulting Langevin dynamics are given by
-$d\mbox{\boldmath$X$}=-(\mathbf{L}+\alpha\mathbf{I})\mbox{\boldmath$X$}dt+\Gamma d\bm{\eta}({t}).$ (5)
-[…]
-$C_{ij}=\frac{\Gamma^{2}}{2}[{\bf L}+\alpha{\bf I}]^{-1}_{ij}.$ (6)
-In other words, from the correlation matrix it is possible to infer the Laplacian, and hence the connectivity of the network.
-We now introduce triadic interactions in the model.
+$\frac{d\mbox{$X$}}{dt}=-(\mathbf{L}+\alpha\mathbf{I})\mbox{$X$}+\Gamma\bm{\eta}({t}).$ (5)
+We remark that the Hamiltonian $\mathcal{H}$ has a minimum for $\mbox{$X$}={\bf 0}$, and its depth increases as the value of $\alpha$ increases.
+In a deterministic version of the model ($\Gamma=0$), the effect of the structural interactions will not be revealed at stationarity.
+However, the Langevin dynamics with $\Gamma>0$ encode the topology on the network.
+Indeed, at equilibrium the correlation matrix $C_{ij}=\mathbb{E}((X_{i}-\mathbb{E}(X_{i}))(X_{j}-\mathbb{E}(X_{j})))$ is given by
+$C_{ij}=\frac{\Gamma^{2}}{2}[{\bf L}+\alpha{\bf I}]^{-1}_{ij},$ (6)
+see Supplementary Information (SI) for details. In other words, from the correlation matrix it is possible to infer the Laplacian, and hence the connectivity of the network.
+We now introduce triadic interactions in the TPM.
 As explained earlier, a triadic interaction occurs when one or more nodes modulate the interaction between another two nodes.
@@
 Namely, we consider the Langevin dynamics
-$d\mbox{\boldmath$X$}=-(\mathbf{L}^{(\text{T})}+\alpha\mathbf{I})\mbox{\boldmath$X$}dt+\Gamma d\bm{\eta}({t}),$ (7)
+$\frac{d\mbox{$X$}}{dt}=-(\mathbf{L}^{(\text{T})}+\alpha\mathbf{I})\mbox{$X$}+\Gamma\bm{\eta}({t}),$ (7)
 obtained from Eq.(2) by substituting the graph Laplacian $\bf L$} with the triadic Laplacian $\mathbf{L}^{(\text{T})}$ whose elements are given by
-$L^{(\text{T})}_{ij}$ =$=\begin{cases}-a_{ij}J_{ij}(\mbox{\boldmath$X$})&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i\neq j,\\[2.5pt] \displaystyle\sum_{k=1}^{N}a_{ik}J_{ik}(\mbox{\boldmath$X$})&\text{if}\leavevmode\nobreak\ \leavevmode\nobreak\ i=j.\end{cases}$ (8)
-Moreover, we assume that the coupling constants $J_{ij}(\mbox{\boldmath$X$})$ are determined by a perceptron-like model that considers all the regulatory nodes of the link $\ell=[i,j]$ and the sign of the regulatory interactions.
-Specifically, if $\sum_{k=1}^{N}K_{\ell k}X_{k}\geq\hat{T}$ then we set $J_{ij}=w^{+}$; if instead $\sum_{k=1}^{N}K_{\ell k}X_{k}<\hat{T}$ then we set $J_{ij}=w^{-}$, with $w_{+},w_{-}\in{}_{+}$ and $w_{+}>w_{-}$.
-Thus, we set
-$J_{ij}(\mbox{\boldmath$X$})$ $=$ $w_{-}+(w_{+}-w_{-})\theta\left(\displaystyle\sum_{k=1}^{N}K_{\ell k}X_{k}-\hat{T}\right),$ (9)
+$L^{(\text{T})}_{ij}$ $=\begin{cases}-a_{ij}J_{ij}(\mbox{$X$})&\text{if}\penalty 10000\ \penalty 10000\ i\neq j,\\[2.5pt] \displaystyle\sum_{k=1}^{N}a_{ik}J_{ik}(\mbox{$X$})&\text{if}\penalty 10000\ \penalty 10000\ i=j.\end{cases}$ (8)
+Moreover, we assume that the coupling constants $J_{ij}(\mbox{$X$})$ are determined by a perceptron-like model that considers all the regulatory nodes of the link $\ell=[i,j]$ and the sign of the regulatory interactions.
+Specifically, if $\sum_{k=1}^{N}K_{\ell k}X_{k}\geq\hat{T}$ then we set $J_{ij}=w^{+}$; if instead $\sum_{k=1}^{N}K_{\ell k}X_{k}<\hat{T}$ then we set $J_{ij}=w^{-}$, with $w_{+},w_{-}\in\real_{+}$ and $w_{+}>w_{-}$.
+Thus,
 […]
-Our model of triadic interactions, therefore, has the crucial property that triadic interactions are higher-order interactions that cannot be reduced to pairs of pairwise interactions.
+[…]
+The TPM for continuous node dynamics in presence of triadic interactions is very general and comprehensively expresses the modulation of structural interactions by other nodes in the network.
+Therefore, the dynamics of TPM cannot be reduced to dynamics exclusively determined by pairwise interactions.
 An important problem that then arises is whether such interactions can be mined from observational data.
-To address this issue, we will develop a new algorithm – that we call the TRIM algorithm – to identify triadic interactions from data, and we will test its performance on the data generated from the model described above.
+To address this issue, we will develop a new algorithm – that we call the TRIM algorithm – to identify triadic interactions from data, and we will test its performance on the data generated from the TPM model described above.
 Figure 2: The TRIM algorithm identifies triples of nodes $X$, $Y$, and $Z$ involved in a putative triadic interaction, starting from the knowledge of the structural network and the dynamical variables associated with its nodes.
-For each putative triple of nodes involved in a triadic interaction (panel (a)) which belong to a network whose structure and dynamics is known (panel (b)), we study the functional behavior of the conditional mutual information $MIz$ (panel (c)), and assess the significance of the observed modulations of $MIz$ with respect to a null model (panel (d)).
+For each putative triple of nodes involved in a triadic interaction (panel (a)) which belong to a network whose structure and dynamics is known (panel (b)), we study the functional behavior of the conditional mutual information MIzMIz (panel (c)), and assess the significance of the observed modulations of MIzMIz with respect to a null model (panel (d)).
 Given a predetermined confidence level, we can use these statistics to identify significant triadic interactions (panel (e)).
@@
 We indicate with $z_{m}$ the quantile of $Z$ corresponding to the percentile $m/P$.
-Therefore, each bin $m_{z}$ indicates data in which $Z$ ranges in the interval $[z_{m},z_{m+1})$.
+Therefore, each bin $m_{z}$ indicates data in which $Z$ ranges in the interval $z_{m},z_{m+1})$.
 We indicate with $\mu(x|z_{m}),\mu(y|z_{m})$ and $\mu(x,y|z_{m})$ the probability density of the variables $X$, $Y$ and the joint probability density of the variables $X$ and $Y$ in each $m_{z}$ bin.
@@
 We thus consider the quantity $MIz(m)=MI(X,Y|Z=z_{m})$ defined as
-[…]
-In order to estimate this quantity, we rely on non-parametric methods based on entropy estimation from $k$-nearest neighbors [41, 42, 43].
+$MIz(m)=\int dx\int dy\mu(x,y|z_{m})\log\left(\frac{\mu(x,y|z_{m})}{\mu(x|z_{m})\mu(y|z_{m})}\right).$
+In order to estimate this quantity, we rely on non-parametric methods based on entropy estimation from $k$-nearest neighbours [41, 42, 43] (see SI for details).
 For each triple of nodes, we visualize the mutual information $MIz$ computed as a function of the $m/P$-th quantiles $z_{m}$ and fit this function with a decision tree comprising $r$ splits.
 In the absence of triadic interactions, we expect $MIz$ to be approximately constant as a function of the $m/P$-th quantiles $z_{m}$, while, in the presence of triadic interactions, we expect this quantity to vary significantly as a function of $z_{m}$.
-The discretized conditional mutual information CMI between $X$ and $Y$ conditioned on $Z$ can be written as
+The discretised conditional mutual information CMI between $X$ and $Y$ conditioned on $Z$ can be written as
 $CMI_{X,Y;Z}=\sum_{m=0}^{P-1}p(z_{m})MIz(m)=\left\langle{MIz}\right\rangle,$ (10)
@@
 This quantity indicates important information about the interaction between the nodes $X$ and $Y$ when combined with the information coming from the mutual information MI given by
-[…]
+$MI_{X,Y}=\int dx\int dy\mu(x,y)\log\left(\frac{\mu(x,y)}{\mu(x)\mu(y)}\right),$
 where $\mu(x),\mu(y)$ are the probability density functions for $X$ and $Y$ and $\mu(x,y)$ is the joint probability density functions of the variables $X$ and $Y$.
-In particular, true interactions between $X$ and $Y$ are typically associated with both high mutual information MI and high conditional mutual information CMI.
-The conditional mutual information CMI, however, is not sensitive to variations in $MIz$ and does not therefore provide the information needed to detect triadic interactions.
-In order to overcome this limitation, we define the following three quantities that measure how much the mutual information between $X$ and $Y$ conditioned on $Z\in[z_{m},z_{m+1})$ changes as $z_{m}$ varies.
+The conditional mutual information, however, is not sensitive to variations in $MIz$ and does not therefore provide the information needed to detect triadic interactions.
+In order to overcome this limitation, we define the following two quantities that measure how much the mutual information between $X$ and $Y$ conditioned on $Z\in[z_{m},z_{m+1})$ changes as $z_{m}$ varies.
 Specifically we consider:
@@
   • (2)
-    the difference $T$ between the maximum and minimum value of $MIz$, given by
+    the difference $T$ between the maximum and average value of $MIz$, given by
     $T=\max_{m=0,\ldots,P-1}\left|MIz(m)-\left\langle{MIz}\right\rangle\right|.$ (12)
-The quantities $\Sigma$, $T$, collectively measure the strength of the triadic interaction under question and can thus be used to mine triadic interactions in synthetic as well as in real data.
+The quantities $\Sigma$ and $T$, collectively measure the strength of the triadic interaction under question and can thus be used to mine triadic interactions in synthetic as well as in real data.
 In order to assess the significance of the putative triadic interaction, we compare the observed values of these variables to the results obtained with given null models.
@@
 Note that if we consider $\mathcal{N}$ realizations of the null model, we cannot estimate probabilities smaller than $1/\mathcal{N}$.
-Therefore, if in our null model we observe no value of $\Sigma^{\textup{ran}}$ larger than the true data $\Sigma$, we set the conservative estimate $p_{\Sigma}=1/\mathcal{P}$.
+Therefore, if in our null model we observe no value of $\Sigma^{\textup{ran}}$ larger than the true data $\Sigma$, we set the conservative estimate $p_{\Sigma}=1/\mathcal{N}$.
 A similar procedure is applied also to $p_{T}$.
-In this work we have considered two types of null models.
+To assess this significance we consider two types of null models.
 The first is the randomization null model obtained by shuffling the $Z$ values, to give $\mathcal{N}$ randomization of the data, i.e.
@@
 The second is the maximum likelihood Gaussian null model between the three nodes involved in the triple $X$,$Y$,$Z$.
-We note that the adoption of these two null models allows us to also potentially capture non-monotonic relationships between $MIz$ and $z$, thereby going beyond underlying monotonic assumptions made in [46].
+Specifically, the Gaussian null model uses the mean and covariance of the timeseries of $X,Y$ and $Z$ to define a multivariate normal distribution from which samples are randomly drawn, thereby providing surrogate timeseries values for the considered triple.
+We note that the use of these two null models allows us to identify also non-monotonic relationships between $MIz$ and $z$, thereby going beyond underlying monotonic assumptions made elsewhere [46].
 The first null model disrupts the temporal correlations between the timeseries of the node $Z$ and the timeseries the two nodes $X$ and $Y$ at the endpoints of the considered edges.
 Therefore this null model is robust with respect to the presence of possible outliers in the dataset.
-However this first null model may overlook confounding network effects that affect correlations between the dynamical variables.
+However, this first null model may overlook confounding network effects that affect correlations between the dynamical variables.
 The second null model more efficiently captures correlations between the dynamical state of the three considered nodes due to network effects but is more sensitive to the presence of outliers in the data.
-To increase confidence, we therefore combine the insights coming from both null models (see SI for details).
+To increase confidence, we therefore combined the insights coming from both these null models (see SI for details).
 For each triple, the function $MIz(m)$ is fitted with a decision tree with two splits.
 In this way, three different intervals of values of $Z$ are identified, each corresponding to a distinct functional behavior of the correlation functions between the variables $X$ and $Y$.
-While the method would allow for more than two splits of the decision three in general, here we have chosen three splits as this is the minimum number of splits that will allow us to capture non-trivial functional behaviour of $MIz$, like non-monotonicity.
-In practice this will be the best choice in presence of limited data and a small number of bins as in the Gene Expression data we will analyse in the following.
-We further characterize significant triples by examining their entropic score function $S$, which is used to characterize their corresponding functional behaviour, as described in the SI.
-As we will discuss below, the algorithm performs well on our paradigmatic model.
-Moreover, while the results obtained with the randomized surrogate data might neglect potentially important network effects, this shortcoming of this approach is mitigated by the cross-validation with the Gaussian null model.
+While our method in principle allows for more than two splits of the decision tree, for illustrative purposes we have chosen two splits since this is the minimum number of splits needed to capture non-trivial functional behavior in $MIz$, such as non-monotonicity.
+In practice this choice of two splits will also be the best choice when data is limited, such as the gene expression data we will analyze in the following section.
+We also further characterize significant triples by examining their normalized entropic score function $S\in[0,1]$, which is used to characterize their corresponding functional behavior.
+Specifically, th entropic score $S$ classifies the diversity of each of the joint distribution functions of $X$ and $Y$ conditioned on $Z$ for each interval obtained through the decision tree (see SI for details).
+As we will discuss below, the algorithm performs well on data obtained from the TPM.
+In this case, we also observe that true triadic triples are characterized by a high entropic score $S$.
+On real data, the results obtained with the TRIM algorithm using randomized surrogate data might neglect potentially important network effects, this shortcoming is mitigated by performing an additional validation using the Gaussian null model and the entropic score $S$ (see SI for the full pipeline of TRIM).
+Figure 3: We consider a network with $N=10$ nodes, $L=12$ edges, and $\hat{L}=5$ triadic interactions (panel (a)).
+Panels (b) and (c) display the effect of triadic interactions on the Mutual Information profile MIzMIz.
+Panel (b) shows MIzMIz for the triple $[4,9,5]$ involved in a positive triadic interaction.
+Panel (c) shows MIzMIz for the triple $[1,2,6]$ that is not involved in a triadic interaction.
+In all panels simulations were run to $t_{\textup{max}}=4,000$ with a timestep of $dt=10^{-2}$.
+For the analysis we consider $40,000$ time steps.
+The parameters of the model are: $\alpha=0.05,\hat{T}=10^{-3},\Gamma=10^{-2}$, $w^{+}=8,w^{-}=0.5$, number of bins $P=400$.
 Figure 4: Representative results for triples of nodes involved in triadic interactions in the continuous model with triadic interactions.
+Results for the triple $[4,9,5]$ of the network in Figure 3 of the main text, which is triadic, are shown.
 The joint distributions of variables $X$ and $Y$ conditional on the values of $Z$ are shown in panel (a).
-Panel (b) shows the behavior of $MIz$ as a function of the values of $z_{m}$, which clearly departs from the constant behavior expected in absence of triadic interactions.
+Panel (b) shows the behavior of MIzMIz as a function of the values of $z_{m}$, which clearly departs from the constant behavior expected in absence of triadic interactions.
 Panel (c) presents the decision tree for fitting the $MIz$ functional behavior and determining the range of values of $Z$ for which the most significant differences among the joint distributions of the variables $X$ and $Y$ conditioned on $Z$ are observed.
-For more details about the simulation results, see the Supplementary Information (SI).
-Figure 4: We consider a network with $N=10$, nodes $L=12$ edges, and $\hat{L}=5$ regulatory interactions (panel (a)).
+The parameters used are the same as in Figure 3.
+Figure 5: We consider the network in Figure 3(a).
 The time series obtained by integrating the stochastic dynamics of the proposed dynamical model for triadic interactions (Eq.(7)) are analyzed with the TRIM algorithm.
-Panel (b) displays the Receiver Operating Characteristic curve (ROC curve) obtained by running TRIM with $P=400$ bins and $\mathcal{N}=10^{3}$ realizations of the null model on these synthetic time series, using $\Theta_{\Sigma}$ to score for different parameters values indicated in the legend.
-The time series is simulated up to a maximum time $t_{max}=4000$ with a $dt=10^{-2}$.
-For the analysis, we consider $40000$ time steps (see the SI for details).
-The parameters of the model are: $\hat{T}=10^{-3}$, $w^{+}=8$, $w^{-}=0.5$, and $\alpha$ and $\Gamma$ as indicated in the figure legend.
-IV.2 Validation of the TRIM algorithm on the triadic interaction model
-To demonstrate the efficacy of the TRIM algorithm we first considered a single network (see Figure 4$(a)$) of $N=10$ nodes, $L=12$ edges and $\hat{L}=5$ triadic interactions (each formed by a single node regulating a single edge) on top of which we consider the triadic dynamical model proposed in Sec.
+Panel (a) displays the Receiver Operating Characteristic curve (ROC curve) obtained by running TRIM with $P=400$ bins and $\mathcal{N}=10^{3}$ realizations of the null model on these synthetic time series, using $\Theta_{\Sigma}$ to score for different parameters values indicated in the legend.
+Panel (b) displays the corresponding Precision-Recall curve (PR curve) obtained by running TRIM with the same parameters.
+The timeseries are simulated up to a maximum time $t_{max}=4000$ with a $dt=10^{-2}$.
+For the analysis, we consider $40,000$ time steps (see the SI for details).
+The parameter of the model are: $\hat{T}=10^{-3}$, $w^{+}=8$, $w^{-}=0.5$, and $\alpha$ and $\Gamma$ as indicated in the figure legend.
+IV.2 Validation of the TRIM algorithm on the triadic perceptron model
+In order to discuss the phenomenology of the TPM we first considered a representative network (see Figure 3) of $N=10$ nodes, $L=12$ edges and $\hat{L}=5$ triadic interactions (each formed by a single node regulating a single edge) on top of which we consider the TPM proposed in Sec.
 III.
-Simulations of Eq.
-(7) on this network show a strong dependence of $MIz(m)$ on $m$ for the triples of nodes involved in triadic interactions, with greater significance for smaller values of $\alpha$.
-Figure 3 shows the joint distributions $P_{\alpha}(X,Y)$ of $X$ and $Y$ for each interval $\alpha$ of values $Z$ determined by the decision tree.
+We found that data obtained from the TPM on this network shows a strong dependence of $MIz(m_{z})$ on $m_{z}$ for the triples of nodes involved in triadic interactions, with greater significance for smaller values of $\alpha$.
+Figure 3 shows the difference between the $MIz(m)$ profile of a triple that is involved in a triadic interaction compared to a triple that is not, demonstrating how triadic interactions modulate the $MIz(m_{z})$ profile.
+Moreover, Figure 4 shows, for a given triadic interaction involving nodes X, Y and Z, the joint distributions $\mu_{\delta}(X,Y)$ of $X$ and $Y$ for each interval $\delta$ of values $Z$ determined by the decision tree.
 The results provide evidence of this interesting dynamical behavior of the triadic model in the case of a positive regulatory interaction.
-Note that our analysis of the function $MIz(m)$ also allows us to distinguish between positive and negative regulatory interactions.
-These are associated respectively by an increase in $MIz$ for larger values of $z_{m}$ or decrease of $MIz$ In the Supplementary Figures S1-S2, we display further examples of the function $MIz$ for triadic triples.
-We observe that as the parameter $\Gamma$ is raised, i.e.
-the noise increases, which is reflected in the increased variability of the $MIz$ functional behavior.
-In order to provide an overall assessment of the performance of TRIM in mining triadic interactions on this network, in Figure 4$(b)$ we display the Receiver Operating Characteristic curve (ROC curve) based on the score $\Theta_{\Sigma}$, measured with respect to the first null model which we found to perform well for a wide range of parameters.
-As it is apparent from the ROC curve, the TRIM algorithm performs better for higher values of $\alpha$ and smaller values of $\Gamma$ (see Figure 4$(b)$).
-For any values of the parameters, we notice that the false positives are more likely to involve short-range triples, i.e.
-triples in which the regulator node $Z$ has short (small) structural network distance with the two nodes $X$ and $Y$.
-Figure 5: Performance of the TRIM algorithm on a benchmark network with triadic interactions.
-Each data point shows measures for a triple of nodes $X$, $Y$ and $Z$, namely $\Theta_{\Sigma}$ and CMI, the proposed $\Sigma$-score and conditional mutual information between $X$ and $Y$, respectively.
-The colour of each point corresponds to the value of $S$, which characterises the entropy of the triple.
-The synthetic data comes from a network with $100$ nodes, $110$ edges, and $25$ triadic interactions.
-All possible triples are represented in the scatter plot.
-Stars are the true triadic triples.
-We conclude our discussion by testing the TRIM algorithm on a much larger model network, to demonstrate its scalability.
-To this end, we consider a random network of $100$ nodes, $110$ edges, and $25$ triadic interactions (see SI for details).
-In Figure 5 we report the results of the TRIM algorithm for all the triples involving all edges, $14$ of which come from the pool of edges with triadic interactions.
-For each edge, all possible triples are analysed while for each edges we retain only the $5$ most significant triples according to $\Theta_{\Sigma}$.
+Note that the analysis of the form of the function $MIz(m_{z})$ also allows us to distinguish between positive and negative regulatory interactions, which are associated with an increase or a decrease in $MIz$ for larger values of $m_{z}$ respectively.
+In the Supplementary Figures S1-S2, we display further examples of the function $MIz$ for triadic triples.
+We observe the increased variability of the $MIz$ functional behavior as the parameter $\Gamma$ is raised, i.e.
+the noise increases.
+These results confirm the main general principle on which the TRIM algorithm is based, i.e.
+that the conditional mutual information $MIz$ is modulated by triadic interactions.
+To make this observation precise, we examined the performance of the TRIM algorithm in mining triadic interactions from synthetic data.
+We first considered the network shown in Figure 3, and using the score $\Theta_{\Sigma}$, we evaluated the Receiver Operating Characteristic (ROC) curve and Precision Recall (PR) curve for different values of the dynamical parameters (see Figure 5).
+Both the ROC curve and the PR curve (which addresses the limitations of the ROC curve for imbalanced datasets) indicate that the TRIM algorithm performs well on data produced by the TPM, with a better performance for higher values of $\alpha$.
+For all parameter values, we noticed that false positives are more likely to involve short-range triples, i.e.
+triples in which the regulator node $Z$ is close (in the structural network) to the end-points $X$ and $Y$ of the target edge.
+Figure 6: Performance of the TRIM algorithm on a random network with triadic interactions.
+(a) Each data point represents a given triple of nodes $X$, $Y$ and $Z$.
+The $y$-axis shows $\Theta_{\Sigma}$, while the $x$-axis shows the CMI between $X$ and $Y.$ The colour of each point corresponds to the value of $S$, which characterizes the entropic score of the triple.
+The synthetic data comes from a structural random Erdös-Renyi network with $100$ nodes, and average degree $c=4$, to which $25$ triadic interactions between random edges and random nodes have been added.
+For the modelling of the network we used $\alpha=0.06,\Gamma=1.4\times 10^{-2},t_{max}=1500,w_{+}=18,w_{-}=0.2$ and for the analysis with TRIM we looked at $3000$ data points and $P=100$ bins.
+We display top 5 triples for each edge according to $\Theta_{\Sigma}$ that are below our $p$ threshold for the randomization null model.
+The Triples below are represented in the scatter plot and they all display an entropic score $S>0.5$.
+Stars are the true triadic triples which are characterized by high $\Theta_{\Sigma}$.
+Crosses are the triples that can be excluded by performing TRIM with the Gaussian Null model.
+(b) Histogram of the $\Theta_{\Sigma}$-values for all the triples of the network (in light blue), and for the triples corresponding to the $25$ true triadic interactions only (in dark blue).
+(c) Histogram of the $\Theta_{\Sigma}$-values observed in a network of the same topology and with the same dynamical parameters for which all the triadic interactions have been removed (orange).
+Figure 7: Maximal Spanning Tree of the relevant genes from the gene expression data.
+Edges and their edge weights were obtained from the Protein-Protein interaction network.
+Red-coloured nodes are nodes with biological significance, that is that play a critical role in AML [47, 48, 49].
+Node size is proportional to the node degree.
+Figure 8: Application of the TRIM algorithm to gene expression data.
+Panel (a) shows the results of TRIM for the significant triples in the AML dataset.
+The scatter plot shows $\Theta_{\Sigma}$ ($y$-axis) versus the CMI ($x$-axis).
+The colour of each point corresponds to the value of its entropic score $S$.
+Here we display only those triples with pp-value 0.001 or less in the randomization null model and that have not been excluded by the Gaussian null model (for details about these triples see SI).
+Circles are triples whose links all appear in the minimum spanning tree, and squares indicate triples involving genes with biological relevance.
+Panels (b)-(c) display the conditional distributions for two example triples: both are identified by the TRIM algorithm with high significance, suggesting a meaningful biological association.
+Panel (b) shows the triple $X$ = GATA1, $Y$ = KLF1, $Z$ = ETV1.
+According to the randomized surrogate null model, this triadic interaction has $p_{\Sigma}$-value 0.001, $\Theta_{\Sigma}$ = 4.75, $\Sigma$ = 0.44, $S=0.64$; panel (c) shows the results for the triple $X$ = HOXB3, $Y$ = MEIS1, $Z$ = GLIS3 involving two biologically relevant genes.
+According to the randomized surrogate null model, it has $\Theta_{\Sigma}$ = 3.98, $p_{\Sigma}$ = 0.001, $\Sigma$ = 0.38, $S=0.60$.
+These results indicate that the TRIM algorithm is effective in identifying triadic interactions in a small network generated using the TPM.
+To examine the scalability of this methodology we also tested the TRIM algorithm on a much larger model network.
+To this end, we consider a random Erdös-Renyi network of $100$ nodes, and average degree $c=4$ to which we added $25$ random triadic interactions (i.e., between randomly chosen nodes to randomly chosen edges), imposing the condition that each edge is at most regulated by a single node for simplicity.
+The results of this analysis are shown in Figure $6(a)$, in which we provide statistics for all possible node triples in the network (the majority of which are not triadic interactions).
+For each edge, we retained only the $5$ most significant triples according to $\Theta_{\Sigma}$.
 By conditioning on the value of third node, for each of these connected nodes we also record the conditional mutual information CMI.
@@
 Stars indicate triples that are involved in a triadic interactions (see SI for details).
-The true triadic interactions are indeed found for triples with high $\Theta_{\Sigma}$ score, while CMI and $S$ span between high and intermediate values.
-Figure 6: Maximal Spanning Tree of the relevant genes from the gene expression data.
-Edges and their edge weights are obtained from the Protein-Protein interaction network.
-Red-coloured nodes are nodes with biological significance, that is that play a critical role in AML[47][48][49].
-Node size is proportional to the node degree.
-Figure 7: (A) Scatter plot which displays $\Theta_{\Sigma}$ (Y-axis) and the CMI (X-axis) value between two genes, denoted $X$ and $Y$ , conditioned by a third gene, denoted $Z$.
-The colour corresponds to the value of our entropy measure $S$.
-Here only triples are depicted that rank among the 300 highest for CMI and MI and have $p$-value 0.001 or less in the randomization null model.
-Circles are triples resulting from the MST, whereas squares are triples involving genes with biological relevance.
-Crosses are triples that have $p$-value biggger than $0.0005$ with the Gaussian version null model.
-(B) The results of the decision tree allow us to distinguish three intervals in the value of Z in which the conditional distribution of the gene expression of $X$ and $Y$ differs the most.
-Plots of these conditional distributions are provided for two exemplary triples: both display a high significance score for the triadic interaction, indicating a meaningful biological association.
-The top row shows three biologically relevant genes, (X = HOXB3, Y = MEIS1, Z = GLIS3, $\Theta_{\Sigma}$ = 5.2, $p_{\Sigma}$ = 0.0001, $\Sigma$ = 0.38); the bottom row shows a triples identified by the MST.
-(C) has $p_{\Sigma}$-value 0.0001 in the random version.
-(X = GATA1, Y = KLF1, Z = ETV1, $\Theta_{\Sigma}$ = 6.1, $\Sigma$ = 0.44)
+True triadic interactions are found for triples with high $\Theta_{\Sigma}$ while CMI span between high and intermediate values.
+This result confirms the very good performance of the TRIM algorithm on the data coming form the TPM.
+To test the statistical robustness of the TRIM algorithm we also conducted the same analysis (i.e., on the same structural network with the same dynamical parameters) in which all triadic interactions were removed.
+The results of this analysis are shown in Figure 6(c).
+In this case, and as expected, the TRIM algorithm did not identify any statistically significant triadic interactions.
+This analysis indicates that the TRIM algorithm is able to identify true triadic interactions with a low false positive discovery rate (compare Figures 6(b)-(c)).
 V Detecting triadic interactions in gene-expression data
-Searching for trigenic interactions in gene-expression is a problem of major interest in biology.
-For instance, understanding the extent to which a modulator promotes or inhibits the interplay between a transcription factor and its target gene is crucial for deciphering gene regulation mechanisms.
-In order to address this question with our method, we considered gene-expression data from Acute Myeloid Leukemia (AML), extracted from the Grand Gene Regulatory Network Database [50, 51].
-Exhaustive mining of all potential triadic interactions among each triple of nodes in the AML dataset is computationally very demanding (it would require testing of $>260M$ triplets).
-Instead, we here focus on edges associated with known biophysical interactions, as identified in the human Protein-Protein Interaction network (PPI)  [50].
-To do so, we considered the subgraph of the human PPI network that contains all the genes/proteins included in the AML gene expression data.
-The number of edges in this subgraph is $M=42,511$.
-We first selected a set of triples involving genes known to be associated to AML with gene X and Y connected by an edge in the PPI (see SI for details).
-Moreover we selected additional triples according to a criteria based on the positions of the three genes in the Maximum Spanning Tree (MST) which only includes 621 edges (see Figure 6).
-For each edge in the MST, connecting gene X with gene Y we consider all genes Z within a distance of 4 from both the X and Y as candidate regulatory nodes, i.e.
+Searching for triadic interactions in gene-expression is a problem of major interest in biology.
+For instance, understanding the extent to which a modulator promotes or inhibits the interplay between a transcription factor and its target gene is crucial for deciphering gene regulation mechanisms [29].
+In order to address this question with our method, we considered a gene-expression dataset associated with Acute Myeloid Leukemia (AML), extracted from the Grand Gene Regulatory Network Database [50, 51].
+Exhaustive mining of all potential triadic interactions from every putative triple of nodes in the AML dataset is computationally very demanding (it would require testing of >260M>260M triples) and likely, due to the sheer number of triples being tested, to result in false positives and/or interactions of less biological importance.
+Moreover, such a brute-force approach would not account for other sources of important biological information, such as putative interactions derived from other experimental sources.
+To account for such information, we therefore focused our analysis on edges between nodes associated with known biophysical interactions, as identified in the human Protein-Protein Interaction network (PPI) [50].
+To do this, we considered the connected subgraph of the human PPI network that contains all the genes/proteins included in the AML gene expression data and their associated edges This network, which contains $622$ nodes and $42,511$ edges, formed the structural network for our analysis [52].
+To start, we focused on triples involving genes known to be associated with AML, in which the end-points $X$ and $Y$ of the target are directly connected in the PPI network (see SI for details).
+We then selected additional triples according to their positions in the PPI network’s Maximum Spanning Tree (MST), which only includes 621 edges (see Figure 7).
+In order to focus on triples for which network effects are likely to be less pronounced, for each edge in the MST, connecting gene/protein $X$ with gene/protein $Y$ we considered all genes $Z$ within a distance of 4 from both the $X$ and $Y$ as candidate regulatory nodes, i.e.
 the third node in the triple (see SI for details).
-For each considered triple of genes we use $\Theta_{\Sigma}$ as the significance score, with $P=5$ bins, using $\mathcal{N}=5\times 10^{3}$ realizations of the randomization null model.
-These results have strong correlations with the results obtained considering $\Theta_{T}$ as the significance score (see Supplementary Figures S3-S4 for comparison).
-Figure 7 shows the result of the TRIM algorithm for a selected set of the analysed triples with $p_{\Sigma}<0.001$.
-Specifically, the figure displays each triple in the plane $\Theta_{\Sigma}$ and CMI with a colour code encoding the value $S$.
+For each considered triple of genes we assessed its significance using $\Theta_{\Sigma}$ as the significance score, with $P=5$ bins, using $\mathcal{N}=5\times 10^{3}$ realizations of the randomization null model (very similar results were obtained using $\Theta_{T}$ as the significance score, see Supplementary Figures S3-S4 for a comparison).
+Figure 8 shows the results of the TRIM algorithm for those triples with $p_{\Sigma}<0.001$.
 Note that for each selected edge only the top $5$ triples ranked according to the $\Theta_{\Sigma}$ score are depicted.
 Squares indicate triples chosen from biologically relevant genes for AML.
-Crosses indicate triples deemed insignificant in comparison with the Gaussian null model, Figures 7B and 7C provide examples of conditional distributions of two exemplary triples with high entropic score $S$.
-Both triples show evidence of triadic interactions: the triple in panel B is a member of the MST, while C is an example of a triple chosen from known biologically relevant genes for AML.
-For further examples of results obtained on more gene-expression triples, see the SI for details.
+The triples deemed insignificant according to the Gaussian null model, are not shown here.
+The interested reader can see their visualization in Figure $S6$ of the SI.
+Figure 8 provide examples of conditional distributions of two illustrative triples that rank high in TRIM.
+Both triples show evidence of a triadic interaction: the triple in panel (b) is a member of the MST, while (c) is an example of a triple chosen from known biologically relevant genes for AML.
+Further example triples are shown in the SI.
 Interestingly, among the significant triples, we detected also triples in which the modulation of the mutual information is non-monotonic (see SI for details).
-Many of the genes involved in the $50$ highest ranking triples have already been linked to AML in the literature (see SI for Table 2 for a list of highly significant triples and Figure S6 with links to the literature associating the involved genes with AML).
-Interestingly, genes such as PAX4, SOX14 and VAX2 which appear in multiple triples have been found to be over- or under-expressed in patients with AML.
-Remarkably, 2 such genes, ASCL1 and NEUROG2, are involved in $5$ triples, suggesting that they may play a particularly important part in disease progression [52].
-In total, $61\%$ of genes in the triples have a known association with AML.
+Many of the genes involved in the $50$ highest ranking triples have already been linked to AML in the literature (see SI for Table S3 for a list of highly significant triples and Table S4 with links to the literature associating the involved genes with AML).
+In total, $84\%$ of the top 50 Triples include at least one gene that has a known association with AML.
 VI Conclusions
-This work provides a comprehensive information theory-based framework to model and mine triadic interactions directly from dynamic observations and for previous knowledge of the underlying pairwise network.
-The generative model we propose demonstrates that the presence of a triadic interaction leads to the modulation of the mutual information between the two end nodes of an edge (X and Y) by a third regulatory node Z.
-Via this model we show that to detect triadic interactions it is necessary to go beyond standard pairwise measures, such as the mutual information.
-Moreover, also standard higher-order statistical measures such as the conditional mutual information that only account for the average effect of the third regulatory node Z on the mutual information between X and Y falls short in characterizing triadic interactions.
-Our proposed approach, implemented in the TRIM algorithm, mines triadic interactions based on the quantification of the statistically significant modulation of the mutual information between the two linked nodes conditioned on the third regulator node.
-This algorithm is tested and validated on our generative model using paradigmatic random networks with triadic interactions and used to mine putative triadic interactions from gene expression data.
-From the network theory point of view, this work opens new perspectives in the active field of modelling and inference of higher-order interactions and can be extended in many different directions, for instance by exploring the effect of triadic interactions on the dynamical state of nodes associated with discrete variables.
+This work provides a comprehensive information theory-based framework to model and mine triadic interactions directly from dynamic observations.
+The TPM we propose demonstrates that the presence of a triadic interaction leads to systematic variations in the mutual information between the two end nodes of the edge involved ($X$ and $Y$).
+Via this model we have shown that to detect triadic interactions it is necessary to go beyond standard pairwise measures, such as the mutual information.
+Importantly, standard higher-order statistical measures, such as the conditional mutual information, which accounts for the average effect of the third regulatory node $Z$ on the mutual information between the target nodes $X$ and $Y$ are also insufficient to identify triadic interactions.
+Our proposed approach, implemented in the TRIM algorithm, mines triadic interactions by identifying statistically significant variations in the mutual information between the two linked nodes conditioned on the third regulator node.
+To demonstrate the efficacy of this algorithm we have tested and validated it on a new dynamical model (that we denote the TPM) and shown how it can identify triadic interactions in randomly generated triadic interaction networks.
+We also used it to mine putative triadic interactions from gene expression data, and connect the putative interactions with meaningful biology.
+From the network theory point of view, this work opens new perspectives in the active field of modelling and inference of higher-order interactions and can be extended in many different directions, for instance by exploring the effect of triadic interactions on the dynamical state of nodes associated with discrete variables or including time delays in the regulation.
 From the biological point of view, our results may inspire further information-theoretic approaches to genetic regulatory network inference.
@@
 Code availability
-The Python package triaction is available on GitHub at the following link: https://github.com/anthbapt/triaction
+The Python package TRIM is available on GitHub at the following link: https://github.com/anthbapt/TRIM
 Acknowledgments
-This research utilized Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT, https://doi.org/10.5281/zenodo.438045.
-M.N.
-A.B., R.
-S.-G.
-and G.
-B.
-acknowledge support from the Turing-Roche North Star partnership.
+This research utilized Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT, https://doi.org/10.5281/zenodo.438045. This work was sponsored by the Turing-Roche Strategic Partnership.
+SUPPLEMENTAL MATERIAL
+Appendix A Supplementary information on the TRIM algorithm
+A.1 Estimation of the Mutual Information
+The Mutual Information $MIz(m)$ between variables $X$ and $Y$ conditioned on the value $z_{m}$ of the variable $Z$ (putative regulator node of the triadic interaction) is calculated using k-nearest-neighbour entropy estimation to be able to estimate it only by using the time series values.
+[41, 42, 43], $MIz(m)$ is defined as:
+$MIz(m)=\int dx\int dy\mu(x,y|z_{m})\log\left(\frac{\mu(x,y|z_{m})}{\mu(x|z_{m})\mu(y|z_{m})}\right).$ (S-1)
+By assuming some metric, neighbours of a chosen point can be ranked by distance.
+The idea is to estimate from the average distance to the nearest neighbour, averaging over all.
+This can be used to estimate the logarithms of the entropies using the Kozachenko-Leonenko-Estimator [41].
+A.2 Entropic score $S$ for significant triples
+In order to identify and classify the significant triples $[X,Y,Z]$ involving node $X$ and $Y$ whose interaction is modulated by node $Z$, we introduce an entropic score function $S$ which characterizes how diverse the conditional joint distributions ${\mu_{\delta}}(X,Y)$ of $X$ and $Y$ conditioned on $Z$ in each of the obtained intervals $\delta\in\{1,2,3\}$ are.
+Dividing the plane $X,Y$ in ${P}^{2}$ squares $(i,j)$ (by binning $X$ and $Y$ in P{P} bins each) with $n^{(\delta)}_{ij}$ data points, we can calculate the participation ratio $Y_{2}^{(\delta)}$ [53, 54]
+$Y_{2}^{(\delta)}=\sum_{i=1}^{{P}}\sum_{j=1}^{{P}}\left(\frac{n_{ij}^{(\delta)}}{\mathcal{N}^{(\delta)}}\right)^{2},$ (S-2)
+where $\mathcal{N}^{(\delta)}=\sum_{i=1}^{{P}}\sum_{j=1}^{{P}}{n_{ij}^{(\delta)}}$.
+The inverse of the partition function is known to measure the effective number of square bins in which the distribution is localized.
+We can then introduce the normalized entropic score $S$ as
+$S=-{\frac{1}{3\ln P^{2}}}\sum_{\delta=1}^{3}\ln Y_{2}^{(\delta)}.$ (S-3)
+The entropy $S$ is low if all the conditional distributions ${\mu_{\delta}}(X,Y)$ are very localized while it acquires large values if all the conditional distributions are delocalized.
+We adopt a threshold $S=0.5$ in order to retain triples with $S>0.5$ indicating that in average the conditional distributions associated to these triples have more than $\sqrt{P^{2}}$ significantly populated bins.
+A.3 Pipeline of TRIM
+In order to select for the significant triples, we combine information coming from the two considered null models (the surrogate randomised data and the maximum likelihood Gaussian model) and the entropic score $S$.
+First we select the set of triples of interest, and we choose which observable to consider, either $\Sigma$ or $T$.
+Note that we have found that the results obtained considering either one of the two observables are highly correlated (see section on gene-expression results).
+With respect to the randomized null model, the Gaussian null model more efficiently captures correlations between the dynamical state of the three considered nodes due to network effects.
+However, the second null model is more sensitive to the presence of outliers in the data but it is still focusing only on the three nodes in the triple and neglects more collective network effects.
+In order to further reduce the set of relevant triples, we screen out triples with low entropic score.
+Thus, we define the following TRIM pipeline:
+  • (a)
+    For each link of interest between variables X and Y consider all the triples X,Y, Z where Z is any given possible regulator node (any node different from X and Y) and calculate $\Theta_{\Sigma}$ or $\Theta_{T}$ together with their corresponding $p$-values.
+Select only triples with p-values smaller than a threshold (typically taken $5\times 10^{-3}$ or $1\times 10^{-3}$).
+  • (b)
+    Screen out all the triples which have a high p-value according to the second (Gaussian) null model.
+Here a high p-value indicates a p-value higher than a threshold typically taken the same as in point (a)).
+  • (c)
+    Rank the remaining triples that rank high according to the $\Theta$ score calculated according to the randomized surrogate data.
+For each link between node X and Y select only the top 5 triples [X,Y,Z].
+  • (d)
+    From the remaining triples screen out all the triples which have a low entropic score $S$ (for instance $S<0.5$).
+Appendix B Supplementary information on the Triadic Perceptron Model (TPM)
+B.1 Derivation of Eq.(6) of the main text
+Let us consider the stochastic dynamics Eq.(5) in the main text, driving the TPM in the absence of triadic interactions, i.e.
+$\frac{d{\bf X}}{dt}=-({\bf L}+\alpha{\bf I}){\bf X}+\Gamma\bm{\eta}({t})$ (S-4)
+where $\mathbf{L}$ is the graph Laplacian of the networks $\alpha,\Gamma\in\mathbb{R}^{+}$ are two real parameters of the model, and $\bm{\eta}$ is a Gaussian white noise with
+$\left\langle{\eta_{i}(t)}\right\rangle=0,\quad\left\langle{\eta_{i}(t)\eta_{j}(t^{\prime})}\right\rangle=\delta_{ij}\delta(t-t^{\prime}).$ (S-5)
+Here we want to derive Eq.(6) of the main text expressing the correlation matrix $C_{ij}=\mathbb{E}((X_{i}-\mathbb{E}(X_{i}))(X_{j}-\mathbb{E}(X_{j})))$ at equilibrium, i.e.
+$C_{ij}=\frac{\Gamma^{2}}{2}[{\bf L}+\alpha{\bf I}]^{-1}_{ij}.$ (S-6)
+[…]
+$\frac{\partial P({\bf X})}{\partial t}=-\sum_{i=1}^{N}\frac{\partial}{\partial X_{i}}[({\bf L}+\alpha{\bf I}){\bf X}]_{i}P({\bf X})+\frac{\Gamma^{2}}{2}\sum_{i,j=1}^{N}\frac{\partial^{2}}{\partial^{2}X_{i}}P({\bf X}).$ (S-7)
+At stationarity, by imposing ${\partial P({\bf X})}/{\partial t}=0$ and solving for $P({\bf X})$, we obtain a multivariate Gaussian distribution
+$P({\bf X})=\mathcal{C}\exp\left[-\frac{1}{2}{\bf X}^{\top}{\bf C}^{-1}{\bf X}\right],$ (S-8)
+where $\mathcal{C}$ is the normalization constant, and $\bf C$} indicates the covariance matrix is given by Eq.(S-6).
+Appendix C Supplementary Information on the TPM dynamical behaviour
+Here we provide supplementary information on the dynamics of the TPM (Eq.
+S-4) on the network with $N=10$ nodes, $L=12$ edges and $\hat{L}=5$ regulatory interactions shown in Figure 4 of the main text.
+We consider the time series obtained by integrating the stochastic dynamics of the proposed dynamical model for triadic interactions within continuous variables.
+The time series is simulated up to a maximum time $t_{\textup{max}}=4000$ with a $dt=10^{-2}$ leading to $4\times 10^{5}$ data points.
+For our analysis we consider the last $200,000$ data points sampled every fifth point leading to $40,000$ time steps.
+In the Supplementary Figures S-1 and S-2 we report the analysis conducted on the triadic triple [4,9,5] already considered in Figure 4 but for different parameter values: for Figure S-1 we have $\alpha=0.01,\Gamma=10^{-2}$, for Figure S-2 we have $\alpha=0.05,\Gamma=5\times 10^{-2}$.
+The other parameter values are: $\hat{T}=10^{-3},w^{+}=8,w^{-}=0.5$, number of bins $P=400$.
+Figure S-1: Exemplary results obtained for a triple of nodes involved in a triadic interaction in the continuous model with triadic interactions.
+The joint distributions of variables $X$ and $Y$ conditional on the values of $Z$ is shown in panel (a).
+Panel (b) displays the functional behaviour of MIzMIz as a function of the values of $z_{m}$ which clearly departs from the constant behaviour expected in absence of triadic interactions.
+Panel (c) presents the decision tree for fitting the $MIz$ functional behaviour and determining the range of values of $Z$ for which the most significant differences among the joint distributions of the variables $X$ and $Y$ conditional to $Z$ are observed.
+The time series is simulated up to maximum time $t_{\textup{max}}=4000$ with $dt=10^{-2}$.
+For the analysis we consider $40,000$ time steps.
+The parameters of the model are: $\alpha=0.01,\hat{T}=10^{-3},$ $\Gamma=10^{-2}$, $w^{+}=8,w^{-}=0.5$, number of bins $P=400$.
+The analysis is done for the triple [4,9,5], of the network in Figure 3 of the main text, which is triadic.
+Figure S-2: Exemplary results obtained for a triple of nodes involved in a triadic interaction in the continuous model with triadic interactions.
+The joint distributions of variables $X$ and $Y$ conditional on the values of $Z$ is shown in panel (a).
+Panel (b) displays the functional behaviour of MIzMIz as a function of the values of $z_{m}$ which clearly departs from the constant behaviour expected in absence of triadic interactions.
+Panel (c) presents the decision tree for fitting the $MIz$ functional behaviour and determining the range of values of $Z$ for which the most significant differences among the joint distributions of the variables $X$ and $Y$ conditional to $Z$ are observed.
+The time series is simulated up to maximum time $t_{\textup{max}}=4000$ with $dt=10^{-2}$.
+For the analysis we consider $40,000$ time steps.
+The parameters of the model are: $\alpha=0.05,\hat{T}=10^{-3},$ $\Gamma=5\times 10^{-2}$, $w^{+}=8,w^{-}=0.5$, number of bins $P=400$.The analysis is done for the triple [4,9,5], of the network in Figure 3 of the main text, which is triadic.
+Appendix D Supplementary information on analysis of gene-expression dataset
+D.1 Methods
+We choose the gene-expression associated with Acute Myeloid Leukemia (AML) extracted from the Grand Gene Regulatory Network Database [50, 51] for our TRIM analysis.
+Testing all possible combinations of triples of genes is computationally too demanding.
+To still make the analysis rigorous, we focus on the most important genes and the genes we suspect are most likely to be involved in triadic interactions.
+Hence, we utilize the Protein-Protein Interaction (PPI) network associated with AML, which captures digenic processes, to choose edges forming the triples in our study.
+This selection is motivated by studies indicating that for gene expression data, most of the genes involved in trigenic processes are also involved in digenic processes [35].
+We consider all the genes in the AML datasets that are connected in the PPI network.
+Each edge in the PPI network has an assigned edge weight.
+This allows us to calculate the maximum spanning tree (MST) (Figure 7).
+The MST has diameter $16$ and includes $622$ nodes and $621$ edges.
+‘EP300’ is the node with the largest node degree.
+To further reduce the number of triples that are being analysed, only triples where the distance of potential regulator node and edge are at least $4$ in the maximal spanning tree are being chosen.
+This is to reduce the effect of the triple being a triangle where all nodes are connected by an edge and to filter out dependencies from nodes of the triple being connected trough only a small number of edges.
+Using this strategy $343,194$ triples are being analysed.
+To further decrease computational time, the number of realisations that the $\Sigma$ or $T$ values are being compared to is adjusted based on the confidence interval for a triadic interaction of the triple being investigated.
+First $10$ realisations are run.
+If the $p_{\Sigma}$-value is bigger than $0.3$, the triple is being categorised as non-triadic.
+In the other case, 100 realisations will be run and if the $p_{\Sigma}$-value of the new result is smaller than 0.05, 1000 realisations will be run.
+This is to quickly exclude any triplets that have extreme low probability of being triadic and only focus on the triples where there is uncertainty whether they are triadic or not.
+Our analysis of the data showed that several genes had an expression profile dominated by outliers that can strongly affect our analysis (see for an example the gene-expression profile of gene TGIF2LX in Figure S-3).
+For each gene $i$ we define an outlier score $O_{i}$ given by
+$O_{i}=\max_{\alpha}\frac{|e_{i\alpha}-\left\langle{e_{i}}\right\rangle|}{\sqrt{\left\langle{e_{i}^{2}}\right\rangle-\left\langle{e_{i}}\right\rangle^{2}}}$ (S-9)
+where $e_{i\alpha}$ is the gene-expression of gene $i$ in sample/patient $\alpha$, $\left\langle{e_{i}^{k}}\right\rangle=\sum_{\alpha}e_{i\alpha}^{k}/Q$ where $Q$ is the number of samples.
+The maximum z-score $O_{i}$ defined in Eq.(S-9) is a good measure to detect outliers.
+We thus remove the top $50$ genes that score highest according to the $O$-score, corresponding to a cut-off $O\simeq 6$ as can be retrieved from a Figure S-4.
+These are the outliers for which triples containing them are removed: [TGIF2LX, DBX2, GSX1, POU4F2, EVX2, EN2, ONECUT1, PAX2, POU1F1, MNX1, HOXC10, BARX2, HOXD12, HOXC12, NKX2-5, FOXC2, HOXC11, PROX1, HMX2, FOXF1, HOXC8, NR0B1, PAX1, GSC, DMRT3, HOXD13, POU3F3, FOXA2, HOXD9, HNF4G, FOXF2, SP8, HMX3, TLX1, FOXD1, FOXA1, PHOX2B, SOX21, DMRT1, FOXQ1, VSX1, HOXC6, EGR4, NKX2-1, ZIC2, LBX1, RXRG, EMX2, SPZ1 ].
+Having removed the outlier genes from our analysis, we test for triadic interactions using the randomisation null model while the Gaussian null model, being more sensitive to the outliers is only used to exclude triples that are not significant.
+Figure S-3: Samples of gene TGIF2LX: Strength of the expression of the gene plotted against each individual patient tested.
+The biggest outlier is at position $17$ with a value of $2.097$, whereas the remaining values are smaller than $0.65$.
+Figure S-4: Cut-off for the top $50$ genes with high outlier values. The black vertical line indicates the cut-off.
+For each edge, we select only the five triples with the highest $\Theta_{\Sigma}$-value because these will be the most likely candidates for a triadic interaction.
+This is consistent with what we find for the synthetic data where triples containing the same edge can have higher $\Theta_{\Sigma}$ or $p_{\Sigma}$ values but only the highest ranking is the true triadic triple.
+To verify that these are triadic and the dependencies do not stem from edges connecting the regulatory node with the other two nodes in the triple, we repeat the analysis for a subset of these triples with the Gaussian version of the null model realisations.
+D.2 Supplementary results
+The analysis reported in Figure 8 of the main-body of the paper focuses on the results obtained by using the $\Sigma$ statistics.
+As supplementary material, in Figure S-5 we report the decision trees corresponding to the conditional distribution of the triples shown in Figure 8 and in Figure S-6 we report the triples screened out according to the TRIM pipeline because they have a $p$-value greater than $0.001$ according to the Gaussian null model.
+Figure S-5: Decision trees belonging to the conditional distribution of the triples shown in Figure 8. Entropy for the top triple is $S=0.64$, entropy for the bottom triple is $S=0.60$.
+Figure S-6: The role of the Gaussian null model in screening out triples in the TRIM pipeline applied to the AML gene expression data.
+The scatter plot displays $\Theta_{\Sigma}$ (Y-axis) verse the CMI (X-axis).
+The colour corresponds to the value of our entropic score $S$.
+Here only triples that have pp-value 0.001 or less in the randomization null model are shown.
+Crosses (in colour) are triples that have $p$-value higher than 0.001 in the Gaussian null model and therefore they are disregarded from the analysis according to the TRIM pipeline.
+Symbols in gray include the triples that are retained and that figure in Figure 8 of the main body of the paper.
+Specifically these include triples whose links appear in the MST (circles), and triples involving genes with biological relevance (squares).
+In Table S-I we use Kendall’s Tau to compare the ranking obtained using the two $\Theta_{\Sigma}$ and $\Theta_{T}$.
+The Kendall’s Tau is higher than $0.7$ for both random and Gaussian null models so those measures are very similar.
+Table S-I: Comparison of the two measures $\Theta_{\Sigma}$ and $\Theta_{T}$ using Kendall-Tau analysis for the Top 50 Triples in our Gene Expression Analysis.
+statistic $p$-value
+Gaussian: $\Theta_{\Sigma}-\Theta_{T}$ $0.864$ $8.12\times 10^{-19}$
+Random: $\Theta_{\Sigma}-\Theta_{T}$ $0.703$ $5.93\times 10^{-13}$
+As mentioned in the main text the TRIM algorithm can be potentially used to detect triadic interactions with non-monotonic behaviour.
+The joint conditional distributions and the decision tree of one exemplary non-monotonic triples is shown in Figure S-7.
+Figure S-7: An example of non-monotonic triadic interaction.
+The figure (a) display the conditional joined distributions for the triple X = BCL6, Y = PPARD, Z = IRF5, $\Theta_{\Sigma}$ = 4.23, $p_{\Sigma}$ = 0.001, $\Sigma$ = 0.32, $S=0.71$, and (b) the decision tree of the same triple.
+The TRIM algorithm was also used to mine triadic interactions among all the triples containing genes that have been found to be highly relevant for triadic interactions from a biological point of view.
+These genes are HOXA1, HOXA2, HOXA3, HOXA4, HOXA5, HOXA6, HOXA7, HOXA8, HOXA9, HOXB1, HOXB2, HOXB3, HOXB4, HOXB5, HOXB6, HOXB7, HOXB8, HOXB9, PBX3, MEIS1 [47], [48], [49].
+There are $125,244$ triples of this type, 260 have been found to be candidates after applying the TRIM pipeline.
+Especially we observe that the method is able to give a high significance score to the triple HOXA9, HOXA3, MEIS1 that has $p_{\Sigma}$-value $0.001$ and $\Theta_{\Sigma}$ greater than 6 which is a triple of overexpressed genes in AML [47].
+This triple together with the highest ranking biological triples by $\Theta_{\Sigma}$ are shown in Table S-II.
+Moreover in Table S-III we report the data for the top $50$ candidates of triadic interactions obtained using our procedure.
+Finally in Table S-IV we provide the list of the literature on the genes involved on the top scoring triples detected by our analysis of the AML gene-expression data.
+Table S-II: Highest ranking Biological Triples by $\Theta_{\Sigma}$
+Node 1 Node 2 Regulatory Node CMI $\Theta_{\Sigma}$ $\Theta_{T}$ $\Sigma$ $P_{\Sigma}$ $P_{T}$ $S$
+HOXA3 HOXA5 PBX3 0.820316 8.265877 8.048662 0.437625 0.001 0.001 0.511452
+HOXA5 HOXA3 LMX1B 0.973260 7.506890 6.763280 0.401504 0.001 0.001 0.631465
+HOXB5 HOXB6 MEIS1 0.819451 7.205339 6.846796 0.427434 0.001 0.001 0.501928
+HOXA2 HOXA5 MEIS1 0.573735 7.019063 5.996416 0.367280 0.001 0.001 0.596266
+HOXA2 HOXA5 MEF2A 0.704308 6.448953 6.374430 0.357321 0.001 0.001 0.618474
+HOXA3 HOXA2 MEF2A 0.736876 6.229782 5.449352 0.357688 0.001 0.001 0.618665
+HOXA9 HOXA3 MEIS1 0.612700 6.147447 5.840250 0.376088 0.001 0.001 0.516719
+HOXA1 HOXB7 DLX3 0.319020 6.031555 5.996593 0.377072 0.001 0.001 0.719936
+HOXA3 HOXA7 LMX1B 0.796522 5.930516 5.361882 0.315926 0.001 0.001 0.575184
+HOXA9 HOXA7 LMX1B 0.795180 5.921237 6.062873 0.351938 0.001 0.001 0.520730
+HOXB3 HOXB6 HOXB7 0.382231 5.755689 4.911239 0.320005 0.001 0.001 0.636879
+HOXB6 HOXB3 MAX 0.711938 5.676961 6.039347 0.309040 0.001 0.001 0.535384
+HOXA3 HOXA2 MEIS1 0.632230 5.643140 5.226696 0.343473 0.001 0.001 0.650853
+HOXB1 PBX3 DMRTC2 0.200534 5.582683 5.388206 0.305488 0.001 0.001 0.566194
+HOXB5 HOXB3 HOXA7 0.504192 5.464608 5.008018 0.320871 0.001 0.001 0.519581
+HOXB5 HOXB6 HOXA3 0.719784 5.333075 5.807730 0.357958 0.001 0.001 0.5175458
+HOXA2 HOXA5 LMX1B 0.685202 5.254939 5.041743 0.314368 0.001 0.001 0.710364
+HOXA3 HOXA2 HOXB3 0.614658 5.212560 5.482464 0.325482 0.001 0.001 0.669220
+HOXB2 HOXB3 SPI1 0.992869 5.194840 5.064013 0.338784 0.001 0.001 0.542340
+HOXA1 HOXB8 RUNX3 0.176414 5.148190 4.848085 0.286748 0.001 0.001 0.626724
+HOXA5 HOXA2 SIX3 0.697338 5.129497 5.577772 0.309603 0.001 0.001 0.542347
+HOXB7 HOXB6 MLXIPL 0.511277 5.113759 5.390683 0.361182 0.001 0.001 0.542626
+HOXA7 PBX3 POU4F1 0.555084 5.111917 5.837126 0.312872 0.001 0.001 0.586764
+HOXB8 HOXA1 DLX3 0.206524 5.058267 5.102255 0.278793 0.001 0.001 0.609568
+HOXB6 HOXB7 STAT5B 0.563708 5.048295 4.784504 0.355652 0.001 0.001 0.569296
+HOXA3 HOXA2 HOXB8 0.704156 4.982283 4.596447 0.326155 0.001 0.001 0.5670250
+HOXB1 PBX3 SREBF1 0.238224 4.964400 3.933486 0.290385 0.001 0.001 0.5993946
+MEIS1 HOXA3 NR6A1 0.484028 4.787769 4.564621 0.384765 0.001 0.001 0.5612223
+HOXA3 HOXA2 ERG 0.854079 4.775633 4.719983 0.307792 0.001 0.001 0.5308264
+HOXB3 HOXB5 IRX1 0.681236 4.762596 4.095362 0.287235 0.001 0.001 0.5965832
+HOXB2 HOXB6 HOXB7 0.349210 4.660569 3.497816 0.316489 0.001 0.002 0.6341667
+HOXB5 HOXB6 HOXA9 0.728788 4.625061 3.950498 0.325499 0.001 0.001 0.5515168
+HOXA9 HOXB7 MLXIPL 0.168949 4.618047 3.982716 0.289461 0.001 0.001 0.6315528
+HOXB3 HOXB5 HOXA5 0.547398 4.574438 5.077232 0.286324 0.001 0.001 0.6519540
+HOXB1 HOXA5 MAFA 0.258676 4.553432 4.648329 0.279503 0.001 0.001 0.6090695
+HOXA1 HOXB6 ATF4 0.371474 4.544329 4.315940 0.299968 0.001 0.001 0.6979978
+MEIS1 HOXA5 HEYL 0.488968 4.526637 3.193835 0.382446 0.001 0.002 0.6221348
+HOXB2 HOXB5 IRX1 0.661687 4.473764 4.903444 0.295175 0.001 0.001 0.6057017
+HOXB7 HOXB6 HOXA9 0.346855 4.444519 3.487488 0.328127 0.001 0.002 0.5408434
+HOXB3 HOXB7 CEBPA 0.456098 4.437850 5.265165 0.302595 0.001 0.001 0.7035531
+HOXB8 HOXA1 PLAGL1 0.174812 4.435822 4.227541 0.259337 0.001 0.001 0.6397053
+PBX3 HOXA3 DLX4 0.376259 4.341928 3.383559 0.294993 0.001 0.003 0.5599292
+HOXA1 HOXA5 RARA 0.564074 4.283333 4.080688 0.347871 0.001 0.002 0.7039146
+HOXB7 HOXB5 CEBPA 0.523503 4.248256 4.572174 0.316662 0.001 0.001 0.5829643
+HOXB8 HOXA9 IRX5 0.143996 4.244078 3.743148 0.255930 0.001 0.002 0.5426364
+HOXA7 PBX3 MEIS2 0.481629 4.229047 4.704656 0.291258 0.001 0.001 0.5720851
+HOXA1 HOXA3 JDP2 0.462743 4.227947 4.378695 0.349784 0.001 0.001 0.6529631
+MEIS1 PBX3 FOXI1 0.399453 4.217408 3.765785 0.351739 0.001 0.001 0.5882254
+HOXB7 HOXB6 HOXA2 0.396262 4.213586 3.643030 0.323173 0.001 0.001 0.6443882
+HOXA2 HOXB7 PBX1 0.297987 4.206895 2.752253 0.303415 0.001 0.005 0.6995975
+Table S-III: Triples with the $50$ highest $\Theta_{\Sigma}$-values.
+Chosen from the analysed triples where only the five triples with highest $\Theta_{\Sigma}$-values for each edge are considered and after the removal of $50$ outlier genes.
+Node 1 Node 2 Reg. Node CMI $\Theta_{\Sigma}$ $\Theta_{T}$ $\Sigma$ $P_{\Sigma}$ $P_{T}$ $S$
+GFI1B HMG20B MAFG 0.159214 6.728150 6.919335 0.239888 0.001 0.001 0.656489
+ESR1 ESRRA CUX1 0.136866 6.633394 6.387613 0.257335 0.001 0.001 0.740265
+ATF2 CREB3L2 NKX2-3 0.183120 6.597125 6.372863 0.292662 0.001 0.001 0.699403
+AR ERG DLX3 0.181392 6.419430 6.294301 0.264732 0.001 0.001 0.670266
+GMEB1 GMEB2 GABPA 0.160774 6.393412 6.419077 0.249189 0.001 0.001 0.648156
+TP53 TOPORS RFX1 0.132448 6.245761 6.067981 0.245629 0.001 0.001 0.636249
+SIX3 NR4A3 PPARD 0.126805 6.157518 5.990863 0.226794 0.001 0.001 0.626979
+CREM CREB3 GABPA 0.145051 6.013678 6.087598 0.217777 0.001 0.001 0.705945
+NR1H2 RXRB FOXO1 0.210384 5.997208 5.846367 0.283394 0.001 0.001 0.622814
+AIRE GMEB1 ZSCAN16 0.139642 5.986620 5.997924 0.230049 0.001 0.001 0.742193
+MAFG NFE2 SMARCC1 0.133617 5.963117 5.929599 0.233347 0.001 0.001 0.675628
+MAFG NFE2 MBD2 0.129252 5.962294 5.873812 0.233020 0.001 0.001 0.656413
+EGR2 HOXB2 HOXA9 0.156973 5.865105 5.867066 0.228347 0.001 0.001 0.688156
+MEIS1 PBX1 ZEB1 0.188129 5.842200 5.931817 0.265427 0.001 0.001 0.611296
+WT1 SOX12 ARX 0.172959 5.825481 5.423547 0.310742 0.001 0.001 0.622995
+MYOD1 TCF12 RFX1 0.134786 5.809604 5.792604 0.202536 0.001 0.001 0.574806
+KDM2B ZNF784 TEAD1 0.184460 5.808177 6.087558 0.237961 0.001 0.001 0.699574
+ELF1 TFDP1 KDM2B 0.180421 5.783468 6.248786 0.216104 0.001 0.001 0.665745
+ZNF423 EBF1 ZIC4 0.134866 5.766551 5.550101 0.245472 0.001 0.001 0.714254
+TP53 IRF5 BCL6 0.185771 5.751306 6.407961 0.206465 0.001 0.001 0.612062
+ESR1 XBP1 HEY2 0.217111 5.741680 5.731264 0.262501 0.001 0.001 0.707043
+MEIS3 PBX3 DRGX 0.116614 5.726887 5.518357 0.225603 0.001 0.001 0.640716
+FOS ELK1 RORA 0.128186 5.714594 5.384268 0.256372 0.001 0.001 0.698619
+LMX1B GBX1 MSX2 0.114587 5.677847 5.501699 0.227926 0.001 0.001 0.626511
+CEBPA CEBPB SMARCC1 0.149690 5.675830 5.864768 0.232468 0.001 0.001 0.629392
+AR HOXB13 IRX5 0.127068 5.628973 5.607486 0.227873 0.001 0.001 0.677219
+LEF1 CDX1 CEBPG 0.173025 5.582373 5.740356 0.204470 0.001 0.001 0.589337
+JUN ATF2 IRF6 0.160900 5.540512 5.706251 0.217531 0.001 0.001 0.743458
+LMX1B FEV CEBPE 0.132053 5.459481 5.381255 0.217621 0.001 0.001 0.531703
+DLX6 HAND2 GFI1 0.316684 5.447074 5.452548 0.409997 0.001 0.001 0.5666130
+EGR3 NFATC2 FOSB 0.139154 5.444294 5.407629 0.230926 0.001 0.001 0.7545026
+ARNT2 EPAS1 STAT5B 0.184543 5.421202 5.515499 0.269106 0.001 0.001 0.6476382
+HOMEZ HMBOX1 POU3F2 0.116605 5.417860 5.441199 0.199304 0.001 0.001 0.6313812
+AR RREB1 LMO2 0.114581 5.415712 5.247730 0.221936 0.001 0.001 0.7792968
+NPAS2 RORA TOPORS 0.139477 5.395178 5.542092 0.214479 0.001 0.001 0.7393471
+FOXP2 SP4 TBX3 0.139812 5.377868 5.488782 0.217301 0.001 0.001 0.5843965
+GATA1 IKZF1 DPRX 0.130265 5.357680 5.120758 0.233326 0.001 0.001 0.7042830
+NANOG TET1 PURA 0.142164 5.348978 5.556867 0.212450 0.001 0.001 0.7522146
+KDM2B ZNF740 NOBOX 0.103249 5.343459 5.074403 0.205580 0.001 0.001 0.7296110
+PITX2 LHX3 TBX5 0.156647 5.315290 4.504254 0.193797 0.001 0.001 0.6662439
+AR ERG DDIT3 0.138062 5.307119 5.147755 0.236228 0.001 0.001 0.6881829
+KDM2B ZNF232 MZF1 0.102529 5.306776 5.161613 0.197083 0.001 0.001 0.6727312
+ARID2 SMARCC2 VDR 0.106636 5.273111 5.122535 0.203007 0.001 0.001 0.6993566
+ESR1 NR2C2 HOXD8 0.129780 5.247737 5.008668 0.234352 0.001 0.001 0.7039140
+LMO2 TAL1 ZNF32 0.105264 5.224616 5.238358 0.182179 0.001 0.001 0.7148965
+ARNT EPAS1 REST 0.121260 5.167161 5.191644 0.214579 0.001 0.001 0.7482859
+LMO2 NHLH1 TBX19 0.102305 5.166281 5.103774 0.196432 0.001 0.001 0.7178833
+E2F3 TFDP1 BPTF 0.184846 5.118170 4.863077 0.203638 0.001 0.001 0.6571473
+SMAD4 SMAD5 HOXB6 0.443231 5.113130 5.490316 0.374981 0.001 0.001 0.6213580
+MITF PAX3 MLXIPL 0.131645 5.102654 4.907555 0.173291 0.001 0.001 0.6373616
+Table S-IV: Genes included in the $50$ triples with highest $\Theta_{\Sigma}$-value.
+The list includes genes relevant to AML found in literature (left) and a list of triples (right).
+Genes with relevance to AML have been bolded.
+Gene Reference Gene 1 Gene 2 Regulatory Gene
+ARID2 [55] T. Bluemn, et al. (2022) GFI1B HMG20B MAFG
+ARNT [56] S. Williams, et al. (2010) ESR1 ESRRA CUX1
+BCL6 [57] K. Kawabata, et al. (2021) ATF2 CREB3L2 NKX2-3
+BPTF [58] A. Radzisheuskaya, et al. (2023) AR ERG DLX3
+CEBPA [59] A. Fasan, et al. (2014) GMEB1 GMEB2 GABPA
+CEBPB [60] W. Fu, et al. (2022) TP53 TOPORS RFX1
+CEBPE [61] Li, K. et al. (2019) SIX3 NR4A3 PPARD
+CEBPG [62] Y. Jiang, et al. (2021) CREM CREB3 GABPA
+CREB3 [63] S. Feng, et al. (2020) NR1H2 RXRB FOXO1
+CUX1 [64] M. E. McNerney, et al. (2013) AIRE GMEB1 ZSCAN16
+ELF1 [65] P. Varghese, et al. (2025) MAFG NFE2 SMARCC1
+ELK1 [66] D. Guo, et al. (2023) MAFG NFE2 MBD2
+EPAS1 [67] S. Wang, et al. (2023) EGR2 HOXB2 HOXA9
+ERG [68] G. Marcucci, et al. (2005) MEIS1 PBX1 ZEB1
+ESR1 [69] A. Roma, et al. (2020) WT1 SOX12 ARX
+FEV [70] J. Zhang, et al. (2022) MYOD1 TCF12 RFX1
+FOS [71] F. Yang, et al. (2024) KDM2B ZNF784 TEAD1
+FOSB [72] SH. Luan, et al. (2022) ELF1 TFDP1 KDM2B
+FOXO1 [73] S. Lin, et al. (2014) ZNF423 EBF1 ZIC4
+GATA1 [74] R. Ayala, et al. (2009) TP53 IRF5 BCL6
+GFI1 [75] T. Möröy, et al. (2015) ESR1 XBP1 HEY2
+HOXA9 [76] S. Aryal, et al. (2023) MEIS3 PBX3 DRGX
+HOXB13 [77] Y. Chu, et al. (2018) FOS ELK1 RORA
+HOXB2 [78] O. Lindblad, et al. (2015) LMX1B GBX1 MSX2
+HOXB6 [79] A. Giampaolo, et al. (2002) CEBPA CEBPB SMARCC1
+IKZF1 [80] J. Eckardt, et al. (2023) AR HOXB13 IRX5
+IRX5 [81] S. Nagel, et al. (2022) LEF1 CDX1 CEBPG
+JUN [82] C. Zhou, et al. (2017) JUN ATF2 IRF6
+KDM2B [83] V. van den Boom, et al. (2016) LMX1B FEV CEBPE
+LEF1 [84] K. Feder, et al. (2020) DLX6 HAND2 GFI1
+LMO2 [85] L. Lu, et al. 2023 EGR3 NFATC2 FOSB
+MBD2 [86] K. Zhou, et al. (2021) ARNT2 EPAS1 STAT5B
+MEIS1 [87] U. Thorsteinsdottir, et al. (2001) HOMEZ HMBOX1 POU3F2
+MYOD1 [88] M. Toyota, et al. (2001) AR RREB1 LMO2
+NFATC2 [89] S. D. Patterson, et al. (2021) NPAS2 RORA TOPORS
+NFE2 [90] J. S. Jutzi, et al. (2019) FOXP2 SP4 TBX3
+NKX2-3 [91] S. Nagel, et al. (2021) GATA1 IKZF1 DPRX
+NPAS2 [92] B. Song, et al. (2018) NANOG TET1 PURA
+NR4A3 [93] SC. Lin, et al. (2022) KDM2B ZNF740 NOBOX
+PBX3 [94] G. J. Dickson, et al. (2013) PITX2 LHX3 TBX5
+PURA [95] K. Lezon-Geyda, et al. (2001) AR ERG DDIT3
+RORA [96] C. Snider, et al. (2019) KDM2B ZNF232 MZF1
+SMAD4 [97] Y. Imai, et al. (2001) ARID2 SMARCC2 VDR
+SOX12 [98] H. Wan, et al. (2017) ESR1 NR2C2 HOXD8
+STAT5B [99] B. Maurer, et al. (2019) LMO2 TAL1 ZNF32
+TAL1 [67] Z. Wang, et al. (2023) ARNT EPAS1 REST
+TET1 [100] J. Wang, et al. (2018) LMO2 NHLH1 TBX19
+TP53 [101] K. Barbosa, et al. (2019) E2F3 TFDP1 BPTF
+WT1 [102] R. Rampal, et al. (2016) SMAD4 SMAD5 HOXB6
+ZEB1 [103] W. G. Shousha, et al. (2019) MITF PAX3 MLXIPL